### v1

# PAID: A Framework of Product-Centric Advertising Image Design

Abstract.
In E-commerce platforms, a full advertising image is composed of a background image and marketing taglines. Automatic ad image design reduces human costs and plays a crucial role. For the convenience of users, a novel automatic framework named Product-Centric Advertising Image Design (PAID) is proposed in this work. PAID takes the product foreground image, required taglines, and target size as input and creates an ad image automatically. PAID consists of four sequential stages: prompt generation, layout generation, background image generation, and graphics rendering. Different expert models are trained to conduct these sub-tasks. A visual language model (VLM) based prompt generation model is leveraged to produce a product-matching background prompt. The layout generation model jointly predicts text and image layout according to the background prompt, product, and taglines to achieve the best harmony. An SDXL-based layout-controlled inpainting model is trained to generate an aesthetic background image. Previous ad image design methods take a background image as input and then predict the layout of taglines, which limits the spatial layout due to fixed image content. Innovatively, our PAID adjusts the stages to produce an unrestricted layout. To complete the PAID framework, we created two high-quality datasets, PITA and PIL. Extensive experimental results show that PAID creates more visually pleasing advertising images than previous methods.
Figure 1. Generated advertising images by our methods with only a product subject image, a target size, and taglines as input.
1. Introduction
Advertising images for products play a crucial role in commercial recommendation systems. Visually harmonious, aesthetically pleasing, and attractive advertising images tend to achieve relatively high click-through rates (Wang et al., 2021). With the rapid advancement of generative methods (Goodfellow et al., 2014; Menick and Kalchbrenner, 2019; van den Oord et al., 2016; Ho et al., 2020; Rombach et al., 2022a), it has become possible to automatically generate advertising images by providing a product foreground image, marketing taglines, and target size as input.
To accomplish this task, by referring to existing approaches (Du et al., 2024; Weng et al., 2024; Wang et al., 2024b; Lin et al., 2023; Zhou et al., 2022), we can first utilize heuristics and predefined rules to generate background text prompts. These prompts are then fed into text-to-image inpainting models to generate backgrounds for product foregrounds, resulting in a composite advertisement image. If marketing taglines are required, the layout of these taglines is predicted based on the generated image, and the taglines are rendered onto the image using techniques like attribute prediction. However, this approach suffers from several limitations: a) it cannot adaptively generate background prompts based on the specific content and shape of the product foreground, leading to poor foreground-background harmony and limited diversity; b) the position of the product foreground is fixed and cannot be adjusted according to the descriptions of the product foreground and background prompt, potentially resulting in inappropriate product size and poor composition; and c) when taglines need to be added, the image content is already finalized, restricting the spatial layout available for tagline placement.
Figure 2. Pipeline of PAID. It consists of four stages and generates advertising images centered around product information.
In this paper, we introduce a novel framework called Product-Centric Advertising Image Design (PAID). This framework centers around the integration of product foreground images and marketing taglines to automatically create more harmonious and aesthetic advertising images. Although PAID remains a multistage approach for generating advertising images, we have adjusted the stages and task definitions to optimize existing methods concerning product-background compatibility and the spatial arrangement of product foregrounds and marketing taglines.
As illustrated in Figure 2, we first introduce a task to generate prompts, using a Visual Language Model (VLM) (Dong et al., 2024) to adaptively create foreground and background prompts that better match the scene, placement, and angle based on product foregrounds. Following this, we propose a layout generation task that organizes the arrangement of the graphic elements (marketing taglines and decorative underlays), product foreground, and other nongraphic elements according to the product foreground, background prompt, marketing taglines, and target size. By advancing the layout stage and arranging the product foreground and marketing information concurrently, as opposed to generating the overall image first and then deciding the marketing text layout, the latter becomes less constrained by the image. This allows for a more rational presentation that is suitable for the needs of the content. In this process, the size and position of the product and other nongraphic elements are considered together, which addresses the limitations of adapting the foreground based on prompts in the inpainting task and reduces conflicts between nongraphic elements and marketing information that could adversely affect readability and attractiveness. Next, we train an inpainting text-to-image model based on SDXL (Podell et al., 2024) that can be controlled by input layouts. Combining inpainting, layout, and prompt control effectively is challenging due to potential conflicts among multiple control signals. A LoRA (Hu et al., 2022) adaption strategy is adopted to balance conflicts and facilitate training. Since layout mainly influences higher-level abstract information, we apply layout control only in the deeper blocks of the UNet, combined with a multi-scale training strategy to enhance training stability and efficiency. Finally, we utilize an attribute prediction and graphics rendering module to overlay graphic elements onto the image. To build the framework and verify its effectiveness, we have established a corresponding dataset, collecting a total of 38,017 images with detailed image captions, product foreground images, and layouts of graphic and nongraphic elements.
In summary, our main contributions are as follows:
  • We propose a novel multi-stage framework for advertising image design, which centers on product foregrounds and taglines and enhances the overall content prominence while maintaining the harmony and aesthetics of the images.
  • A VLM is introduced to generate suitable prompts for image generation using the input product foreground image, considering product placement angle and shape.
  • To arrange product foregrounds and taglines more effectively, a VLM-based layout generator is proposed. The generator can mix and arrange graphic and nongraphic elements with the target size and prompt. Accordingly, an image generation model that simultaneously accepts layout, prompt, and foreground control is proposed. Several strategies are introduced to enhance effectiveness and efficiency.
  • We created two datasets: one with 38,017 annotated images for advertising image design and another with 12,247 annotated images for background generation. We use these datasets to train and validate the models in the framework.
2. Related Work
2.1. Advertising Image Design
In recent years, various algorithms are proposed to automate advertising image design in e-commerce scenario (Zhou et al., 2022; Hsu et al., 2023; Li et al., 2023b; Seol et al., 2024; Lin et al., 2023; Li et al., 2023a). Typically, these approaches can be divided into two subcategories: background image-based and foreground-based methods. Background image-based methods adopt a full image as input and design models to predict the layout, followed by attribute prediction and graphics rendering techniques to construct the final ad image. Here, a layout means the class and position of texts, underlays, or other graphic elements. CGL-GAN (Zhou et al., 2022) and DS-GAN (Hsu et al., 2023) design transformer architectures to leverage the content of the background image and its saliency map for generating layout. RADM (Li et al., 2023b) builds a diffusion model and uses the image content and tagline content to derive the layout. It is the first work to consider tagline content. PosterLlama (Seol et al., 2024) leverages the rich design knowledge in Large Language Models (LLMs) to predict the layout. AutoPoster (Lin et al., 2023) introduces a complete pipeline which first generates a layout based on a background image, and then predicts the attributes of graphic elements and renders them onto the image. For these methods, the spatial layout is restricted due to fixed image content, which further restricts the tagline content, such as number and length.
In contrast, P&R (Li et al., 2023a) is a foreground-based method that leverages both the product foreground image and tagline content to predict the layout of the product foreground and taglines. It subsequently designs a module network to outpaint the background and render the texts according to the given background prompt and generated layout. The tagline content is not limited, and the layout is more flexible. However, P&R ignores the harmony between the background prompt and the generated layout. The layout may conflict with background generation. In our PAID framework, the overall layout of graphic elements, product foreground, and other nongraphic elements is generated jointly according to the product foreground, background prompt, and tagline contents, aiming to achieve the best harmony among them.
2.2. Controllable Image Generation
Diffusion models (Ho et al., 2020; Rombach et al., 2022a) have significantly improved image generation quality. To control instances and improve image quality for text-to-image models, some methods introduce layout guidance. These layout-to-image methods can be grouped into training-free and training-based categories. Mostly, training-free layout-to-image methods either manipulate the attention map of instances in prompt (Kim et al., 2023), or compute attention-based loss and back-propagate the gradients in de-noising procedure (Chefer et al., 2023; Xie et al., 2023; Chen et al., 2024). MultiDiffusion (Bar-Tal et al., 2023) conducts inference for each instance and then combines their latent vectors. Training-based methods modify the network architectures and fine-tune the diffusion model with layout-image data (Li et al., 2023c; Wang et al., 2024a; Zhou et al., 2024). GLIGEN (Li et al., 2023c) and InstanceDiffusion (Wang et al., 2024a) add a gated self-attention layer to introduce layout information. MIGC (Zhou et al., 2024) conducts per-instance shading and aggregates all the shaded instances.
For advertising image generation (Du et al., 2024; Wang et al., 2023; Cao et al., 2024; Li et al., 2023a), text-to-image models equipped with inpainting capabilities are frequently utilized. The inpainting technique is employed to preserve product characteristics, while the text prompt is used to describe the background. The foreground image is utilized by incorporating extra channels into the UNet architecture (Rombach et al., 2022a) or employing an inpainting ControlNet (Zhang et al., 2023). To better control the overall layout of advertising images and improve the quality of generated images, we combine inpainting with layout control. Currently, there is limited exploration in this area. SceneBooth (Chai et al., 2025) is the first model to integrate both controls in training, combining gated self-attention layers and inpainting ControlNet to train a model based on SD 1.5 (Rombach et al., 2022b) that can accept layout and foreground controls. However, directly combining the two to train a larger model like SDXL can lead to non-convergence or a decline in the quality of generated images. Therefore, we use several strategies when training the layout-controlled inpainting model to ensure training stability, generation quality, and efficiency.
3. PITA Dataset and PIL Dataset
For automatic product-centric advertising image design, we collect a large-scale Product-Centric Image-Text Advertising (PITA) dataset, gathering 38,017 samples from e-commerce platforms and the CGL dataset (Zhou et al., 2022). 1,000 of them are used for testing. PITA is more diverse than CGL, featuring images with various aspect ratios: 0.684 (51%), 1.0 (33%), 0.667 (14%), and 0.75 (2%). Each image is labeled with prompts (foreground caption and background caption), a product mask, a layout of graphic and nongraphic elements. Each element is represented with a type and a bounding box (bbox). Graphic elements contain “Logo”, “Text”, and “Underlay”. Figure 3 shows some examples in the PITA dataset.
Figure 3. Examples from the PITA dataset, with text, underlay, and logo highlighted in orange, yellow, and red rectangles, respectively.
For layout-controlled background inpainting, we additionally construct a Product-Centric Image Layout (PIL) dataset. The images do not contain graphic elements and labeled with prompts, product masks, and layouts of nongraphic elements. Figure 4 gives some examples. The image scenes cover indoors and outdoors. Product types include furniture, food, clothes, shoes, and so on. PIL contains 12,247 samples in total and 1,000 of them are divided for test.
Figure 4. Examples in PIL dataset.
4. Method
4.1. Overall Framework
As illustrated in Figure 2, our proposed PAID framework consists of four stages: prompt generation, layout generation, background image generation, and graphics rendering.
4.2. Prompt generation
For prompt generation, we fine-tune a large vision-language model $\pi_{prompt}$ based on InternLM-XComposer2-vl (XCP2) (Dong et al., 2024). The method is illustrated in the upper part of Figure 5.
Input data format. For the input of LLM, we first describe the prompt generation task and ask for foreground and background descriptions. For the image input, we crop out the product foreground and feed it into the visual encoder.
Output data format. To enable downstream layout and background image generation tasks, and to better understand the relationship between the foreground and background, we use the prompt generation model $\pi$ to simultaneously predict the foreground description $p_{fore}$ and the background prediction $p_{back}$. We format the output in JSON as it is compatible with the pre-trained VLM and simplifies the following analysis. The output JSON is structured as $p=(p_{fore},p_{back})$.
Training scheme. A standard VLM training scheme involves two stages: pre-training and supervised fine-tuning (SFT) (Liu et al., 2023a). During pre-training, the projector learns to align features between visual and text modalities. In the SFT stage, the projector and LLM are further trained for visual understanding tasks. Since we utilize the VLM trained after the SFT stage, the model can understand input images and follow instructions. We only fine-tune the LLM branch with our dataset to adapt the model to the prompt generation task, using cross-entropy loss as the objective function. The model $\pi_{prompt}$ takes the foreground image $I_{fore}$ and predefined instructions as input, performing two main tasks: generating a foreground description $p_{fore}$ and predicting the background description $p_{back}$. The prompt generation process is as follows:
(1) $p=\pi_{prompt}(I_{fore})$
4.3. Layout generation
In this stage, we introduce the Jointly Predict Graphic and Nongraphic Layout (JPGNL) method, which aims to optimize the arrangement and improve the readability and appeal of advertising images. It predicts the layout of graphic elements, the product, and other nongraphic elements based on the foreground image, prompt, provided taglines, and target size. Similar to prompt generation, we use XCP2 (Dong et al., 2024) for layout generation. The method is shown in the lower part of Figure 5.
Figure 5. The framework of our prompt and layout generation model and data format.
Input data format. For the input to the LLM, we describe the layout generation task and organize input conditions in JSON format. The image input remains the same as in prompt generation. A target canvas aspect ratio is given as condition to support multi-scale layout design. The aspect ratio of the product foreground is provided to help keep its shape. Foreground and background prompts are also given for generating the image layout. For the pre-defined taglines, we list them with empty locations and expect the model to complete them. If a logo is needed, we include it as a condition in the graphic layout section and specify its aspect ratio.
Except for these basic conditions, we also take product characteristics into account. In E-commerce scenarios, some products should not be occluded for better display. While for others, like clothing or human-hold items, partial occlusion is acceptable. To differentiate between these, we introduce a Class-Conditioned Layout Prediction (CCLP) strategy. CCLP provides clear information about the product class and whether foreground occlusion is allowed. For the not-allowing occlusion set (No Occ Set), we describe “The class of subject is [V1]. The bounding boxes of texts should never occlude the subject.” For the allowing occlusion set (Allow Occ Set), the description is “The class of subject is [V2]. The bounding boxes of texts are allowed to occlude the subject.”
Output data format. The output layout is structured in JSON format with three parts: subject layout, other non-graphic layout, and graphic layout. To maintain the shape of the product foreground, we propose a Ratio-Keeping Bbox Representation (RKBR). It is expressed as [x, y, h*r1/r2, h]. (x,y) is the normalized center point. w and h are the normalized width and height. r1 and r2 are the actual aspect ratio of the foreground and canvas respectively. Each value ranges from 0 to 1 with three decimal places. The nongraphic layout contains the location of elements in the background prompt, each having a type (instance name) and bounding box (bbox). The graphic layout defines each element by type, content, and bounding box, covering text, underlay, and logo. The content key is used only for text types to specify its semantic content. Bounding boxes are listed as [x, y, w, h]. The model is expected to supplement underlay adaptively. If logos are required, it is represented in the same way as the product for ratio keeping.
Training scheme. Same as the prompt generation module, we only fine-tune the LLM branch with the visual encoder and projector frozen. The output JSON is structured to predict the layout in a raster scan order, adhering to the ascending order of top-left coordinates. Furthermore, we shuffle the taglines in the input conditional JSON for data augmentation and preventing the model from memorizing the input sequence.
4.4. Background Image generation
In this process, we train an SDXL-based (Podell et al., 2024) image generation model controlled by product foreground, prompt, and layout.
Layout control. Suppose there are $N$ elements to be drawn in an image, represented as $\{(n_{i},l_{i})|i=0,1,...,N-1\}$. We map these into $N$ embeddings as layout input. Here, $n_{i}$ is the element name, and $l_{i}$ is its location. Concretely, we use the text encoder in SDXL to map $n_{i}$ into a global semantic embedding $\textbf{e}_{n,i}$. $l_{i}$ is transferred into normalized top-left and bottom-right coordinates, then fed into a Fourier mapping layer(Tancik et al., 2020), deriving the location embedding $\textbf{e}_{l,i}$. Next, $\textbf{e}_{n,i}$ and $\textbf{e}_{l,i}$ are concatenated and fed into a trainable MLP layer to represent the layout information of the $i$-th element, denoted as $\textbf{e}_{i}$. $\textbf{e}=\{\textbf{e}_{i}|i=0,1,...,N-1\}$ is the layout embeddings. A gated self-attention (GSA) layer (Li et al., 2023c) is inserted between the self- and cross-attention layers in the original UNet to inject layout embeddings. Supposing there are $M$ visual embeddings (V) after the self-attention layer. The output of the GSA layer is
$GSA(\textbf{V},\textbf{E})=\textbf{V}+tanh(\gamma)*SA(cat(\textbf{V},\textbf{E}))[:M]$
, where $\gamma$ is a learnable parameter, $cat$ means concatenation, and $SA$ means self-attention.
Considering that layout is semantic-level information and less related to high-frequency details, we only add GSA layers in the deep blocks of UNet. We refer to this strategy as Deep Layer Control (DLC). Specifically, the layout control is applied in the middle block and the lowest resolution of up blocks. This reduces the parameters and inference latency, and has no negative impact on generation quality. Additionally, we adopt a LoRA Adaptation Training (LAT) strategy that trainable LoRA (Hu et al., 2022) layers are added to UNet, for bridging the gap between the base model and added GSA layers. This further accelerates convergence and increases generation quality. Figure 6 describes how layout control is achieved.
Figure 6. Illustration of layout control.
Training stages. To build the product-reserving layout-to-image model, the training involves two steps. First, we modify the UNet architecture in SDXL by adding the above GSA layers and train it for deriving a layout-to-image model. We use a multi-scale training strategy for faster convergency, first training on 512x512 resolution images and then fine-tuning on 1024x1024 resolution images. Second, we equip the trained model with an inpainting ControlNet and further fine-tune it on 1024x1024 resolution images. This step incorporates the product foreground, achieving full control over the prompt, layout, and product foreground.
4.5. Graphics Rendering
Since there is no effective method to predict the visual attributes of text and other graphic elements, and the graphics rendering module is not the focus of this paper, we design a strategy according to the aesthetic rules to handle graphics rendering. We first sort the text elements by area to decide the font and color, both of which are limited to a specific set (Liu et al., 2024). The selection is based on the similarity to the foreground color and contrast with the background area. Texts grouped by size and position use the same font and color. For the underlay, the color is chosen based on its contrast with the text color and similarity to the background color. The shape is selected from a predefined SVG library based on its size and proportions, with minor adjustments for fitting.
5. Experiments
In this section, we compare our PAID framework and its modules with previous ones. We conducted experiments on the PITA dataset, and the results confirm the effectiveness of our method.
5.1. Implementation Details
For prompt and layout generation, XCP2 (Dong et al., 2024) is used as our backbone. Its visual encoder is built upon CLIP ViT-Large (Radford et al., 2021), processing 490x490 resolution images. The LLM is based on InternLM2 (Cai et al., 2024). We fine-tune the LLM branch on the PITA dataset for one epoch. We train the model using AdamW optimizer with a learning rate of 1e-5. Batch size on each GPU is set to 1 and gradient accumulation of 8 is adopted. The training costs about 14 hours with 8 NVIDIA A100 GPUs. During inference, we use top-p sampling with p set to 0.9 and a sampling temperature of 0.6.
For background image generation, we use SDXL base model as the backbone. During the first step, about 6.8 million high-aesthetic internal images with layout annotations are used. We train the newly added layers with a learning rate of 5e-5 for 20k steps on 512x512 resolution and further fine-tune them with a learning rate of 2e-5 for 20k steps on 1024x1024 resolution. Then in step 2, we introduce an inpainting ControlNet which is pre-trained with millions of advertising images from E-commerce platforms (Creative, 2024), and fine-tune the layout-related layers with a learning rate of 2e-5 for 1k steps, using 12k training data in the collected PIL dataset.
5.2. Evaluation Metrics
Overall pipeline. We evaluate the effectiveness of the overall pipeline from two perspectives: overall visual quality and layout. For visual quality, we employ the Fréchet Inception Distance (FID) (Heusel et al., 2017) and aesthetic score. The aesthetic score is obtained by advertising experts ranking the outputs of different pipelines. The average rank is calculated to represent the mean aesthetic rating, where a lower rank indicates superior creative results. For layout, we adopt the metrics used in previous studies (Zhou et al., 2022; Seol et al., 2024), which are detailed in the subsequent Layout Generation module.
Prompt generation. We evaluated prompt generation quality based on fore-background matching and e-commerce domain score.
  • FBM Rate (Fore-background matching rate): It refers to how well a product foreground fits with a given prompt background. We classify this matching into “reasonable” or “unreasonable” through human annotation, and then determine the proportion of reasonable cases for each method.
  • ED Score (E-commerce domain score): To assess how well the model-generated prompts reflect e-commerce characteristics, we calculate the FID between the CLIP features of these prompts and the ground truth prompts in the test set.
Layout generation. Following previous work (Zhou et al., 2022; Seol et al., 2024), we use graphic metrics and content metrics to evaluate layouts. Graphic metrics focus on relationships between graphic elements, including validity $Val$, alignment $Ali$, overlap $Ove$, and underlay $Und_{l},Und_{s}$. Val means the ratio of elements larger than 0.1% of the canvas. Ali assesses how well elements are spatially aligned. Ove is the average intersection over union (IoU) of all element pairs except for underlays. $Und_{l}$ and $Und_{s}$ checks if underlays enhance non-underlay elements. Content metrics assess the harmony of layout with the background image, including utility $Uti$, occlusion $Occ$, and unreadability $Rea$. Uti evaluates space usage for graphic elements. Occ is the average overlapping area between the graphic elements and products. Rea represents the non-flatness of regions that text elements without underlays are put on. For methods which take taglines as input, we report the text match rate (TMR), to see if the number of generated text bboxes are consistent with input taglines.
Background image generation. We evaluate the quality of generated background images with FID, CLIP-T (Ruiz et al., 2023), CLIP-I (Gal et al., 2023). To verify layout control, we use Grounding-DINO (Liu et al., 2023b) to detect instances and calculate the maximum IoU between detected bboxes and ground-truth. The average of these maximum IoUs is called mIOU. In addition, Grounding-DINO AP scores are also reported.
5.3. Comparison
Compare with automatic pipelines for product advertising image design. Since there are no public automatic methods for advertising image design conditioned on the product foreground, taglines and target size, we design two pipelines for comparison. Pipeline 1 use GPT-4o (OpenAI, 2024) for prompt and layout generation based on the input conditions. We fine-tune the SDXL-based inpainting model on our PIL dataset to obtain the image generation model. The only difference between this method and ours is the absence of layout control and the corresponding training stage. During inference, only the product foreground is controlled by the layout output of GPT-4o. Pipeline 2 first generate a background prompt according to the foreground image using GPT-4o. The image generation model is the same as in Pipeline 1, but during inference, the foreground is placed according to general athetic rules (see details in the Appendix). Next, the image and taglines are processed by the SOTA layout generation model PosterLlama to generate a graphic layout. Last, the results of these two pipelines are processed through the same graphics rendering module as our method to produce the final image. Table 1 shows the quantitative comparison between these pipelines and our PAID framework. PAID outperforms in most metrics, except for slightly higher occlusion compared to Pipeline 1, which places the product centrally with a small bounding box. However, Pipeline 1 uses GPT-4o for layout, lacking product-centric design knowledge, resulting in lower quality. PAID uses trained experts for different tasks in ad design, creating more visually appealing results (see Figure 7). Pipeline 2 fixes the product location, limiting adaptability to product shapes, whereas PAID adjusts layouts based on product shape for better visuals. Additionally, the fixed background of Pipeline 2 can lead to layouts with crowded text boxes, making them hard to read, while our model provides more flexible and readable layouts.
Table 1. Quantitative comparison between automatic pipelines for product-centric advertising image design.
Overall Visual Quality Layout Quality
[…]
Pipeline 1 56.545 2.781 0.0021 0.002 0.8333 0.7872 0.0891 0.0937 0.234
Pipeline 2 43.221 1.814 0.0016 0.0028 0.9994 0.9930 0.0984 0.1158 0.1864
PAID 37.524 1.405 0.0013 0.0017 0.9999 0.9973 0.1367 0.0955 0.1815
Figure 7. Visualization of advertising image designed by different methods.
Compare with prompt generation methods. For prompt comparison, we use GPT-4o and an untrained version of our model (XCP2) as baselines. Both utilize in-context learning (ICL) for optimal performance. The FBM Rate is determined by votes from three annotators for each case. As shown in Table 2, our method outperforms the baselines in both FBM Rate and ED Score. We also use t-SNE for dimensionality reduction to visualize the feature distribution of prompts generated by our model before and after training (see Figure 8). This shows that our model effectively aligns with the e-commerce prompt dataset distribution after training.
Table 2. Evaluation of prompt generation methods
Method FBM Rate $\uparrow$ ED Score $\downarrow$
GPT-4o 97.9% 22.49
XCP2 94.8% 13.04
Finetuned XCP2 98.5% 7.17
Figure 8. T-sne visualization of generated prompts.
Compare with layout generation methods. To verify the advantage of our layout generation model, we compare it with previous background-based methods (DS-GAN (Hsu et al., 2023), RADM (Li et al., 2023b), and PosterLlama (Seol et al., 2024)) and the foreground-based method P&R (Li et al., 2023a). For background-based generation methods, we conduct experiments on two versions of the testing set for fair comparison, named the erased set and the generated set. The erased set contains images with graphic elements erased by LaMa (Suvorov et al., 2022), which may provide hints to put graphic elements on these erased area. The generated set contain images re-generated using their ground-truth prompts (image captions). The generated set has no above hints and conforms to real applications. As for the foreground-based method P&R, since it has not released the codes, we re-implement its core idea: first, generating the layout of the foreground and graphic elements conditioned on the foreground and taglines, and second, generating the final image based on the background prompt, layout, and product foreground. Our method uses the ground-truth prompt, product foreground, and taglines in PITA for prediction. The results are shown in Table 3 and Figure 9. Background-based methods perform relatively poorly in terms of product occlusion and overlap and alignment of graphic elements. This may be due to the fixed background image content, leaving less room for the model flexibility. RADM and DS-GAN, which are insensitive to tagline content, also face issues with generating text boxes that cannot adequately fill the tagline content. The foreground-based method P&R improves in these areas but does not consider the harmony between the product, background prompt, and graphic layout, sometimes leading to unreasonable generated images. Our method, PAID, takes into account the shape of the product foreground, background prompt, and tagline content to decide the overall layout, including graphic and nongraphic elements, resulting in a better overall effect.
Table 3. Quantitative performance comparison of layout generation methods on PITA dataset. Bold number represents the best result in each column except for erased set. BG-based and FG-based are the abbreviations for Background-based and Foreground-based methods, respectively.
[…]
BG-based (Erased) DS-GAN(Hsu et al., 2023) 0.9585 0.0270 0.0058 0.3910 0.0744 0.1816 0.1063 0.1826 -
RADM(Li et al., 2023b) 0.999 0.0411 0.0017 0.9852 0.6934 0.1484 0.0814 0.1722 0.82
PosterLlama(Seol et al., 2024) 0.9984 0.002 0.0026 0.9899 0.9870 0.1226 0.0976 0.1820 0.998
BG-based (Generated) DS-GAN(Hsu et al., 2023) 0.9621 0.0284 0.0080 0.3350 0.0528 0.1701 0.1202 0.2328 -
RADM(Li et al., 2023b) 0.9753 0.0484 0.0201 0.7997 0.2828 0.0657 0.2527 0.2558 0.289
PosterLlama(Seol et al., 2024) 1.0 0.0015 0.0022 0.9990 0.9965 0.1090 0.1208 0.231 0.996
FG-based P&R(Li et al., 2023a) 1.0 0.0012 0.0019 0.9966 0.9929 0.1367 0.1000 0.2008 1.0
Ours 1.0 0.0012 0.0017 0.9976 0.9956 0.1364 0.0973 0.1968 1.0
Figure 9. Qualitative comparison of layout generation methods on PITA test set. Orange: text, Yellow: underlay, Red: logo, Black: Invalid
Compare with layout-controlled inpainting models. We compare our layout-controlled background image generation model with others, including training-based method GLIGEN (Li et al., 2023c) and training-free methods BoxDiff (Xie et al., 2023), TFCLG (Chen et al., 2024), and MultiDiffusion (Bar-Tal et al., 2023). Since these methods lack inpainting ability, we re-implement them with the SDXL-based inpainting ControlNet (Creative, 2024). Quantitative results are reported in Table 4. Training-free methods have weaker spatial control compared to training-based methods. Our trained model outperforms others in FID, mIoU, and AP, indicating better image quality and spatial control. As shown in Figure 10, other models tend to miss some instances when more objects are required, while our model generates them more accurately.
Figure 10. Visualization of images generated by different layout-controlled inpainting models.
Table 4. Quantitative performance comparison of layout-controlled inpainting models on PIL test set.
[…]
BoxDiff(Xie et al., 2023) 29.383 0.315 0.896 0.555 0.037/0.064/0.033
TFCLG(Chen et al., 2024) 29.498 0.315 0.895 0.543 0.033/0.056/0.029
MultiDiffusion(Bar-Tal et al., 2023) 42.359 0.311 0.841 0.5617 0.029/0.052/0.025
GLIGEN(Li et al., 2023c) 27.237 0.312 0.898 0.561 0.047/0.073/0.044
Ours 25.917 0.313 0.906 0.705 0.079/0.127/0.078
5.4. Ablation Studies
Effect of JPGNL in layout generation. Our method predicts the overall layout of graphic and nongraphic elements. Here, we examine the need to predict the layout of nongraphic elements according to the background prompt. As shown in Figure 11, without specifying the position of nongraphic elements, texts placed on complex areas can lead to a messy appearance. By predicting the overall layout, we improve the text readability of advertising images. Additionally, with nongraphic layout predictions, generated images align more closely with background prompts. Without JPGNL, images might miss some objects from the prompts. Besides, in Table 5, we quantitatively analyze the impact of predicting nongraphic layouts on graphic layouts. Predicting nongraphic layouts (w/ JPGNL) can improve metrics related to product occlusion and text readability. However, since it adds complexity to the task, it negatively affects graphic metrics such as element overlap and alignment.
Figure 11. An illustration of the effect of JPGNL. Predicting image layout makes better visual effect.
Table 5. Quantitative ablation study on JPGNL.
[…]
w/o JPGNL 1.0 0.0009 0.0013 0.9995 0.9891 0.1377 0.1001 0.2003
w/ JPGNL 1.0 0.0012 0.0017 0.9976 0.9956 0.1364 0.0973 0.1968
Effect of RKBR and CCLP in layout generation. We validate the effectiveness of the proposed input and output format for layout generation, including RKBR and CCLP. As shown in Table 6, without RKBR, the model may predict a product size which does not match the original aspect ratio about 5% of the time. Note that we regard a difference of less than 1.5% as correctness. Adding CCLP improves the Uti and Occ values on No Occ Set, indicating that CCLP helps the model to distinguish occlusion-allowing sets and others by providing explicit class conditions.
Table 6. Ablation study on input/output format construction of layout generation. O, NO, and AO correspond to the overall set, the No Occ Set, and the Allow Occ Set, respectively. FRC represents Fg Ratio Correctness.
[…]
1.0 0.0012 0.0018 0.9982 0.9971 0.1353/0.1448/0.0876 0.0985/0.0230/0.4258 0.2005 0.955
1.0 0.0012 0.0018 0.9991 0.9912 0.1360/0.1466/0.0816 0.1054/0.0209/0.4491 0.1994 1.0
1.0 0.0008 0.0018 0.9989 0.9906 0.1366/0.1492/0.0851 0.1002/0.0156/0.4282 0.1973 0.957
1.0 0.0012 0.0017 0.9976 0.9956 0.1364/0.1493/0.0849 0.0973/0.0136/0.4316 0.1968 1.0
Effect of DLC and LAT in layout-to-image model. We investigate the effect of deep layer control (DLC) and Lora adaptation training (LAT), with results shown in Table 7. Adding layout control in deep UNet layers performs almost as well as controlling all layers. It reduces parameters and inference costs, which benefits the application. With LoRA adaptation, the model achieves lower FID and higher mIoU, enhancing image quality and spatial control. Parameters and inference costs can be further reduced.
Table 7. Quantitative performance comparison of layout-to-image methods on PIL dataset.
[…]
27.237 0.312 0.898 0.561 0.047/0.073/0.044
27.308 0.312 0.898 0.558 0.045/0.075/0.043
25.675 0.313 0.906 0.696 0.076/0.123/0.075
25.917 0.313 0.906 0.705 0.079/0.127/0.078
6. Conclusion
In this paper, a novel framework named Product-Centric Advertising Image Design (PAID) is proposed to automatically create ad images only based on product foreground images, marketing taglines, and target sizes. PAID consists of four serial stages: prompt generation, layout generation, background image generation, and graphics rendering. First, it generates adaptive prompts according to the semantic and visual information of the product, for the benefit of background image generation. Then, it predicts how graphic and nongraphic elements should be placed according to the prompt, product foreground, and taglines, creating a harmonious overall layout. Next, a layout-controlled inpainting model is utilized for background image generation. A graphics rendering module is then applied to get the final images. We train separate experts to conduct these sub-tasks. Two high-aesthetic and diverse product-centric datasets are created for the convenience of training. Qualitative and quantitative comparisons prove that the proposed PAID framework generates more visually pleasing ad images, compared with previous methods.
Acknowledgements
This work is supported by National Natural Science Foundation of China (NSFC No.62076031 and No.62076036) and Alibaba Research Intern Program.

### v2

@@
 Abstract.
-In E-commerce platforms, a full advertising image is composed of a background image and marketing taglines.
-Automatic ad image design reduces human costs and plays a crucial role.
-For the convenience of users, a novel automatic framework named Product-Centric Advertising Image Design (PAID) is proposed in this work.
-PAID takes the product foreground image, required taglines, and target size as input and creates an ad image automatically.
-PAID consists of four sequential stages: prompt generation, layout generation, background image generation, and graphics rendering.
-Different expert models are trained to conduct these sub-tasks.
-A visual language model (VLM) based prompt generation model is leveraged to produce a product-matching background prompt.
-The layout generation model jointly predicts text and image layout according to the background prompt, product, and taglines to achieve the best harmony.
-An SDXL-based layout-controlled inpainting model is trained to generate an aesthetic background image.
-Previous ad image design methods take a background image as input and then predict the layout of taglines, which limits the spatial layout due to fixed image content.
-Innovatively, our PAID adjusts the stages to produce an unrestricted layout.
-To complete the PAID framework, we created two high-quality datasets, PITA and PIL.
-Extensive experimental results show that PAID creates more visually pleasing advertising images than previous methods.
-Figure 1. Generated advertising images by our methods with only a product subject image, a target size, and taglines as input.
+Creating visually appealing advertising images is often a labor-intensive and time-consuming process.
+Is it possible to automatically generate such images using only basic product information—specifically, a product foreground image, taglines, and a target size?
+Existing methods mainly focus on parts of the problem and fail to provide a comprehensive solution.
+To address this gap, we propose a novel multistage framework called Product-Centric Advertising Image Design (PAID).
+It consists of four sequential stages to highlight product foregrounds and taglines while achieving overall image aesthetics: prompt generation, layout generation, background image generation, and graphics rendering.
+Different expert models are designed and trained for the first three stages: First, we use a visual language model (VLM) to generate background prompts that match the products.
+Next, a VLM-based layout generation model arranges the placement of product foregrounds, graphic elements (taglines and decorative underlays), and various nongraphic elements (objects from the background prompt).
+Following this, we train an SDXL-based image generation model that can simultaneously accept prompts, layouts, and foreground controls.
+To support the PAID framework, we create corresponding datasets with over 50,000 labeled images.
+Extensive experimental results and online A/B tests demonstrate that PAID can produce more visually appealing advertising images.
+Figure 1. Generated advertising images by our methods with a product foreground image, taglines, and a target size as input.
 1. Introduction
-Advertising images for products play a crucial role in commercial recommendation systems.
-Visually harmonious, aesthetically pleasing, and attractive advertising images tend to achieve relatively high click-through rates (Wang et al., 2021).
-With the rapid advancement of generative methods (Goodfellow et al., 2014; Menick and Kalchbrenner, 2019; van den Oord et al., 2016; Ho et al., 2020; Rombach et al., 2022a), it has become possible to automatically generate advertising images by providing a product foreground image, marketing taglines, and target size as input.
-To accomplish this task, by referring to existing approaches (Du et al., 2024; Weng et al., 2024; Wang et al., 2024b; Lin et al., 2023; Zhou et al., 2022), we can first utilize heuristics and predefined rules to generate background text prompts.
-These prompts are then fed into text-to-image inpainting models to generate backgrounds for product foregrounds, resulting in a composite advertisement image.
-If marketing taglines are required, the layout of these taglines is predicted based on the generated image, and the taglines are rendered onto the image using techniques like attribute prediction.
-However, this approach suffers from several limitations: a) it cannot adaptively generate background prompts based on the specific content and shape of the product foreground, leading to poor foreground-background harmony and limited diversity; b) the position of the product foreground is fixed and cannot be adjusted according to the descriptions of the product foreground and background prompt, potentially resulting in inappropriate product size and poor composition; and c) when taglines need to be added, the image content is already finalized, restricting the spatial layout available for tagline placement.
+Advertising images (as shown in Figure 1) are essential for commercial recommendation.
+Visually appealing and attractive images usually have high click-through rates (Wang et al., 2021) but are labor-intensive and time-consuming to create (Vashishtha et al., 2024).
+With the rapid advancement of generative methods (Menick and Kalchbrenner, 2018; Van Den Oord et al., 2016; Ho et al., 2020; Rombach et al., 2021), it is now possible to automatically generate advertising images using only basic product inputs: a product foreground image, marketing taglines, and a target size.
+Currently, generating advertising images using only the basic inputs mentioned above has rarely been explored.
+Existing methods (Du et al., 2024; Weng et al., 2024; Wang et al., 2024b; Lin et al., 2023; Zhou et al., 2022) mainly focus on parts of the problem, such as background or layout generation, rather than offering a complete solution for integrating product foregrounds and taglines.
+Although these methods can be combined, they struggle to effectively highlight product images and taglines while maintaining visual appeal and harmony.
+Specifically, building on existing approaches (Du et al., 2024; Weng et al., 2024; Wang et al., 2024b; Lin et al., 2023; Zhou et al., 2022), we can first utilize predefined rules to create background prompts.
+These prompts are then fed into text-to-image inpainting models to complete backgrounds for product foregrounds.
+Layouts for taglines and decorative elements (collectively referred to as graphic elements) are predicted based on the images and taglines.
+Finally, attribute prediction and graphics rendering techniques are employed to add these graphic elements to the images.
+However, this process has several limitations: a) Lack of adaptability: It cannot generate background prompts that match the specific content and shape of the product, leading to poor foreground-background harmony.
+b) Fixed foreground positioning: The position of the foreground cannot be adjusted based on its characteristics and the background prompt, potentially resulting in inappropriate product size and poor composition.
+c) Restricted tagline placement: When adding taglines, the whole image content is already set, limiting the available space for tagline placement.
 Figure 2. Pipeline of PAID. It consists of four stages and generates advertising images centered around product information.
-In this paper, we introduce a novel framework called Product-Centric Advertising Image Design (PAID).
-This framework centers around the integration of product foreground images and marketing taglines to automatically create more harmonious and aesthetic advertising images.
-Although PAID remains a multistage approach for generating advertising images, we have adjusted the stages and task definitions to optimize existing methods concerning product-background compatibility and the spatial arrangement of product foregrounds and marketing taglines.
-As illustrated in Figure 2, we first introduce a task to generate prompts, using a Visual Language Model (VLM) (Dong et al., 2024) to adaptively create foreground and background prompts that better match the scene, placement, and angle based on product foregrounds.
-Following this, we propose a layout generation task that organizes the arrangement of the graphic elements (marketing taglines and decorative underlays), product foreground, and other nongraphic elements according to the product foreground, background prompt, marketing taglines, and target size.
-By advancing the layout stage and arranging the product foreground and marketing information concurrently, as opposed to generating the overall image first and then deciding the marketing text layout, the latter becomes less constrained by the image.
-This allows for a more rational presentation that is suitable for the needs of the content.
-In this process, the size and position of the product and other nongraphic elements are considered together, which addresses the limitations of adapting the foreground based on prompts in the inpainting task and reduces conflicts between nongraphic elements and marketing information that could adversely affect readability and attractiveness.
-Next, we train an inpainting text-to-image model based on SDXL (Podell et al., 2024) that can be controlled by input layouts.
+In this paper, we propose a novel framework called Product-Centric Advertising Image Design (PAID) to enhance advertising image generation using only basic product inputs.
+While PAID remains a multistage process, we have refined its stages and task definitions to enhance foreground-background compatibility and the spatial arrangement of product foregrounds and taglines.
+As illustrated in Figure 2, we first introduce a task to generate prompts, using a Visual Language Model (VLM) (Dong et al., 2024) to adaptively create background prompts that better match the scene, placement, and angle based on product foregrounds.
+Following this, we propose a layout generation task that arranges the graphic elements (marketing taglines and decorative underlays), product foreground, and other nongraphic elements (objects to be generated from the background prompt).
+This arrangement is based on the product foreground, background prompt, marketing taglines, and target size.
+By arranging the product foreground and marketing information concurrently, as opposed to generating the overall image first and then deciding the graphic layout, it becomes less constrained by the image content and allows for a more rational presentation of the product info.
+Furthermore, considering the nongraphic elements together helps minimize conflicts with taglines that could affect readability and attractiveness.
+Next, we train an inpainting text-to-image model based on SDXL (Podell et al., 2023) that can be controlled by input layouts.
 Combining inpainting, layout, and prompt control effectively is challenging due to potential conflicts among multiple control signals.
-A LoRA (Hu et al., 2022) adaption strategy is adopted to balance conflicts and facilitate training.
-Since layout mainly influences higher-level abstract information, we apply layout control only in the deeper blocks of the UNet, combined with a multi-scale training strategy to enhance training stability and efficiency.
+A LoRA (Hu et al., 2021) adaption strategy is adopted to balance conflicts and facilitate training.
+Since layout mainly influences higher-level abstract information, we apply layout control only in the deeper blocks of the UNet (He et al., 2016), combined with a multi-scale training strategy to enhance training stability and efficiency.
 Finally, we utilize an attribute prediction and graphics rendering module to overlay graphic elements onto the image.
-To build the framework and verify its effectiveness, we have established a corresponding dataset, collecting a total of 38,017 images with detailed image captions, product foreground images, and layouts of graphic and nongraphic elements.
+To build the framework and verify its effectiveness, we have collected and labeled over 50,000 images to establish datasets.
+In testing, PAID demonstrates its ability to produce advertising images with better visual quality.
+It has been deployed on two advertising recommendation scenarios of Taobao, and online A/B tests show that its generated images are more appealing to users.
 In summary, our main contributions are as follows:
-  • We propose a novel multi-stage framework for advertising image design, which centers on product foregrounds and taglines and enhances the overall content prominence while maintaining the harmony and aesthetics of the images.
+  • We propose a novel framework PAID for advertising image design given only a product foreground image, taglines, and a target size.
+It centers on the product foreground and taglines to enhance the overall product prominence and improve the image aesthetics.
   • A VLM is introduced to generate suitable prompts for image generation using the input product foreground image, considering product placement angle and shape.
   • To arrange product foregrounds and taglines more effectively, a VLM-based layout generator is proposed.
-The generator can mix and arrange graphic and nongraphic elements with the target size and prompt.
+It can mix and arrange graphic and nongraphic elements with the target size and prompt.
 Accordingly, an image generation model that simultaneously accepts layout, prompt, and foreground control is proposed.
 Several strategies are introduced to enhance effectiveness and efficiency.
-  • We created two datasets: one with 38,017 annotated images for advertising image design and another with 12,247 annotated images for background generation.
-We use these datasets to train and validate the models in the framework.
+  • We have collected and labeled over 50,000 images to create datasets for advertising image generation.
+The effectiveness of our method has been validated through tests on this dataset and online A/B tests in real advertising scenarios.
 2. Related Work
 2.1. Advertising Image Design
-In recent years, various algorithms are proposed to automate advertising image design in e-commerce scenario (Zhou et al., 2022; Hsu et al., 2023; Li et al., 2023b; Seol et al., 2024; Lin et al., 2023; Li et al., 2023a).
-Typically, these approaches can be divided into two subcategories: background image-based and foreground-based methods.
-Background image-based methods adopt a full image as input and design models to predict the layout, followed by attribute prediction and graphics rendering techniques to construct the final ad image.
-Here, a layout means the class and position of texts, underlays, or other graphic elements.
-CGL-GAN (Zhou et al., 2022) and DS-GAN (Hsu et al., 2023) design transformer architectures to leverage the content of the background image and its saliency map for generating layout.
-RADM (Li et al., 2023b) builds a diffusion model and uses the image content and tagline content to derive the layout.
-It is the first work to consider tagline content.
+In recent years, many methods have been proposed to automate advertising image design (Zhou et al., 2022; Hsu et al., 2023; Li et al., 2023b; Seol et al., 2024; Lin et al., 2023; Li et al., 2023a).
+They can be divided into two subcategories: background-based and foreground-based methods.
+Background-based methods predict graphic layouts based on complete images (not only product foregrounds), followed by attribute prediction and graphics rendering modules to create the final advertising image.
+CGL-GAN (Zhou et al., 2022) and DS-GAN (Hsu et al., 2023) design transformer architectures to leverage image info and generate layouts.
+RADM (Li et al., 2023b) is the first to simultaneously consider the image and tagline content to derive layouts with a diffusion model.
 PosterLlama (Seol et al., 2024) leverages the rich design knowledge in Large Language Models (LLMs) to predict the layout.
-AutoPoster (Lin et al., 2023) introduces a complete pipeline which first generates a layout based on a background image, and then predicts the attributes of graphic elements and renders them onto the image.
-For these methods, the spatial layout is restricted due to fixed image content, which further restricts the tagline content, such as number and length.
+AutoPoster (Lin et al., 2023) introduces a complete pipeline which first generates a layout based on a complete image, and then predicts the attributes of graphic elements and renders them onto the image.
+For these methods, the spatial layout is restricted due to fixed image content, which further restricts the tagline content, including the number and length of taglines.
 In contrast, P&R (Li et al., 2023a) is a foreground-based method that leverages both the product foreground image and tagline content to predict the layout of the product foreground and taglines.
-It subsequently designs a module network to outpaint the background and render the texts according to the given background prompt and generated layout.
+It subsequently designs a module network to outpaint the background and render the taglines according to the given background prompt and generated layout.
 The tagline content is not limited, and the layout is more flexible.
@@
 2.2. Controllable Image Generation
-Diffusion models (Ho et al., 2020; Rombach et al., 2022a) have significantly improved image generation quality.
-To control instances and improve image quality for text-to-image models, some methods introduce layout guidance.
-These layout-to-image methods can be grouped into training-free and training-based categories.
-Mostly, training-free layout-to-image methods either manipulate the attention map of instances in prompt (Kim et al., 2023), or compute attention-based loss and back-propagate the gradients in de-noising procedure (Chefer et al., 2023; Xie et al., 2023; Chen et al., 2024).
-MultiDiffusion (Bar-Tal et al., 2023) conducts inference for each instance and then combines their latent vectors.
-Training-based methods modify the network architectures and fine-tune the diffusion model with layout-image data (Li et al., 2023c; Wang et al., 2024a; Zhou et al., 2024).
-GLIGEN (Li et al., 2023c) and InstanceDiffusion (Wang et al., 2024a) add a gated self-attention layer to introduce layout information.
-MIGC (Zhou et al., 2024) conducts per-instance shading and aggregates all the shaded instances.
+There have been significant developments in text-to-image models (Rombach et al., 2021; Podell et al., 2023).
+To control instances and further improve image quality, some methods introduce layout guidance, dividing into training-free and training-based categories.
+Training-free methods manipulate the attention map (Kim et al., 2023) or compute attention-based loss during de-noising (Chefer et al., 2023; Xie et al., 2023; Chen et al., 2024; Bar-Tal et al., 2023).
+Training-based methods, like GLIGEN (Li et al., 2023c) and InstanceDiffusion (Wang et al., 2024a), modify network architecture, incorporating layout data by adding gated self-attention layers.
 For advertising image generation (Du et al., 2024; Wang et al., 2023; Cao et al., 2024; Li et al., 2023a), text-to-image models equipped with inpainting capabilities are frequently utilized.
-The inpainting technique is employed to preserve product characteristics, while the text prompt is used to describe the background.
-The foreground image is utilized by incorporating extra channels into the UNet architecture (Rombach et al., 2022a) or employing an inpainting ControlNet (Zhang et al., 2023).
+The inpainting technique is employed to preserve product characteristics, while the prompt is used to describe the background.
+The foreground image is utilized by incorporating extra channels into the UNet architecture (Rombach et al., 2021) or employing an inpainting ControlNet (Zhang et al., 2023).
 To better control the overall layout of advertising images and improve the quality of generated images, we combine inpainting with layout control.
 Currently, there is limited exploration in this area.
-SceneBooth (Chai et al., 2025) is the first model to integrate both controls in training, combining gated self-attention layers and inpainting ControlNet to train a model based on SD 1.5 (Rombach et al., 2022b) that can accept layout and foreground controls.
+SceneBooth (Chai et al., 2025) is the first model to integrate both controls in training, combining gated self-attention layers and inpainting ControlNet to train a model based on SD 1.5 (Rombach et al., 2021) that can accept layout and foreground controls.
 However, directly combining the two to train a larger model like SDXL can lead to non-convergence or a decline in the quality of generated images.
@@
 3. PITA Dataset and PIL Dataset
-For automatic product-centric advertising image design, we collect a large-scale Product-Centric Image-Text Advertising (PITA) dataset, gathering 38,017 samples from e-commerce platforms and the CGL dataset (Zhou et al., 2022).
+For automatic image design of product-centric advertising, we collect a large-scale Product-Centric Image-Tagline Advertising (PITA) dataset, gathering 38,017 samples from e-commerce platforms and the CGL dataset (Zhou et al., 2022).
 1,000 of them are used for testing.
-PITA is more diverse than CGL, featuring images with various aspect ratios: 0.684 (51%), 1.0 (33%), 0.667 (14%), and 0.75 (2%).
-Each image is labeled with prompts (foreground caption and background caption), a product mask, a layout of graphic and nongraphic elements.
+PITA is a dataset of varying sizes, featuring images with four distinct aspect ratios: 0.684, 1.0, 0.667, and 0.75.
+Each image is labeled with prompts (foreground caption and background caption), a product mask, and a layout of graphic and nongraphic elements.
 Each element is represented with a type and a bounding box (bbox).
-Graphic elements contain “Logo”, “Text”, and “Underlay”.
-Figure 3 shows some examples in the PITA dataset.
-Figure 3. Examples from the PITA dataset, with text, underlay, and logo highlighted in orange, yellow, and red rectangles, respectively.
+The graphic elements contain “Logo”, “Tagline”, and “underlay”.
+Some examples of data visualization are in the appendix.
 For layout-controlled background inpainting, we additionally construct a Product-Centric Image Layout (PIL) dataset.
-The images do not contain graphic elements and labeled with prompts, product masks, and layouts of nongraphic elements.
-Figure 4 gives some examples.
-The image scenes cover indoors and outdoors.
-Product types include furniture, food, clothes, shoes, and so on.
+The images do not contain graphic elements and are labeled with prompts, product masks, and layouts of nongraphic elements.
 PIL contains 12,247 samples in total and 1,000 of them are divided for test.
-Figure 4. Examples in PIL dataset.
 4. Method
@@
 Training scheme.
-A standard VLM training scheme involves two stages: pre-training and supervised fine-tuning (SFT) (Liu et al., 2023a).
+A standard VLM training scheme involves two stages: pre-training and supervised fine-tuning (SFT) (Liu et al., 2024a).
 During pre-training, the projector learns to align features between visual and text modalities.
@@
 The prompt generation process is as follows:
-(1) $p=\pi_{prompt}(I_{fore})$
+(1) $p=\pi_{prompt}(I_{fore}).$
 4.3. Layout generation
-In this stage, we introduce the Jointly Predict Graphic and Nongraphic Layout (JPGNL) method, which aims to optimize the arrangement and improve the readability and appeal of advertising images.
-It predicts the layout of graphic elements, the product, and other nongraphic elements based on the foreground image, prompt, provided taglines, and target size.
+In this stage, we introduce the Jointly Predict Graphic and Nongraphic Layout (JPGNL) method, which optimizes the arrangement and improves the tagline readability and image appeal.
+It predicts the layout of graphic elements, the product, and other nongraphic elements based on the foreground image, prompt, taglines, and target size.
 Similar to prompt generation, we use XCP2 (Dong et al., 2024) for layout generation.
@@
 The image input remains the same as in prompt generation.
-A target canvas aspect ratio is given as condition to support multi-scale layout design.
+A target canvas aspect ratio is given as a condition to support multi-scale layout design.
 The aspect ratio of the product foreground is provided to help keep its shape.
@@
 For the not-allowing occlusion set (No Occ Set), we describe “The class of subject is [V1].
-The bounding boxes of texts should never occlude the subject.” For the allowing occlusion set (Allow Occ Set), the description is “The class of subject is [V2].
-The bounding boxes of texts are allowed to occlude the subject.”
+The bounding boxes of taglines should never occlude the subject”.
+For the allowing occlusion set (Allow Occ Set), the description is “The class of subject is [V2].
+The bounding boxes of taglines are allowed to occlude the subject”.
+Figure 4. Layout control in the image generation model.
 Output data format.
-The output layout is structured in JSON format with three parts: subject layout, other non-graphic layout, and graphic layout.
+The output layout is structured in JSON format with three parts: subject layout, other nongraphic layout, and graphic layout.
 To maintain the shape of the product foreground, we propose a Ratio-Keeping Bbox Representation (RKBR).
@@
 w and h are the normalized width and height.
-r1 and r2 are the actual aspect ratio of the foreground and canvas respectively.
+r1 and r2 are the actual aspect ratio of the foreground and canvas, respectively.
 Each value ranges from 0 to 1 with three decimal places.
 The nongraphic layout contains the location of elements in the background prompt, each having a type (instance name) and bounding box (bbox).
-The graphic layout defines each element by type, content, and bounding box, covering text, underlay, and logo.
-The content key is used only for text types to specify its semantic content.
+The graphic layout defines each element by type, content, and bounding box, covering tagline, underlay, and logo.
+The content key is used only for tagline types to specify its semantic content.
 Bounding boxes are listed as [x, y, w, h].
@@
 The output JSON is structured to predict the layout in a raster scan order, adhering to the ascending order of top-left coordinates.
-Furthermore, we shuffle the taglines in the input conditional JSON for data augmentation and preventing the model from memorizing the input sequence.
+Furthermore, we shuffle the taglines in the input conditional JSON for data augmentation and prevent the model from memorizing the input sequence.
 4.4. Background Image generation
-In this process, we train an SDXL-based (Podell et al., 2024) image generation model controlled by product foreground, prompt, and layout.
+In this process, we train an SDXL-based (Podell et al., 2023) image generation model controlled by product foreground, prompt, and layout.
 Layout control.
-Suppose there are $N$ elements to be drawn in an image, represented as $\{(n_{i},l_{i})|i=0,1,...,N-1\}$.
+Suppose that there are $N$ elements to be drawn in an image, represented as $\{(n_{i},l_{i})|i=0,1,...,N-1\}$.
 We map these into $N$ embeddings as layout input.
@@
 The output of the GSA layer is
-$GSA(\textbf{V},\textbf{E})=\textbf{V}+tanh(\gamma)*SA(cat(\textbf{V},\textbf{E}))[:M]$
-, where $\gamma$ is a learnable parameter, $cat$ means concatenation, and $SA$ means self-attention.
-Considering that layout is semantic-level information and less related to high-frequency details, we only add GSA layers in the deep blocks of UNet.
+$GSA(\textbf{V},\textbf{E})=\textbf{V}+tanh(\gamma)*SA(cat(\textbf{V},\textbf{E}))[:M]$,
+where $\gamma$ is a learnable parameter, $cat$ means concatenation, and $SA$ means self-attention.
+Considering that the layout is semantic-level information and less related to high-frequency details, we only add GSA layers in the deep blocks of UNet.
 We refer to this strategy as Deep Layer Control (DLC).
@@
 This reduces the parameters and inference latency, and has no negative impact on generation quality.
-Additionally, we adopt a LoRA Adaptation Training (LAT) strategy that trainable LoRA (Hu et al., 2022) layers are added to UNet, for bridging the gap between the base model and added GSA layers.
+Additionally, we adopt a LoRA Adaptation Training (LAT) strategy in which trainable LoRA (Hu et al., 2021) layers are added to UNet, to bridge the gap between the base model and added GSA layers.
 This further accelerates convergence and increases generation quality.
 Figure 4 describes how layout control is achieved.
-Figure 6. Illustration of layout control.
 Training stages.
-To build the product-reserving layout-to-image model, the training involves two steps.
-First, we modify the UNet architecture in SDXL by adding the above GSA layers and train it for deriving a layout-to-image model.
+To build the product reservation layout-to-image model, training involves two steps.
+First, we modify the UNet architecture in SDXL by adding the above GSA layers and train it to derive a layout-to-image model.
 We use a multi-scale training strategy for faster convergency, first training on 512x512 resolution images and then fine-tuning on 1024x1024 resolution images.
@@
 4.5. Graphics Rendering
-Since there is no effective method to predict the visual attributes of text and other graphic elements, and the graphics rendering module is not the focus of this paper, we design a strategy according to the aesthetic rules to handle graphics rendering.
-We first sort the text elements by area to decide the font and color, both of which are limited to a specific set (Liu et al., 2024).
+Since there is no effective method to predict the visual attributes of graphic elements, and the graphics rendering module is not the focus of this paper, we design a strategy according to the aesthetic rules to handle graphics rendering.
+We first sort the tagline elements by area to decide the font and color, both of which are limited to a specific set (Liu et al., 2024b).
 The selection is based on the similarity to the foreground color and contrast with the background area.
-Texts grouped by size and position use the same font and color.
-For the underlay, the color is chosen based on its contrast with the text color and similarity to the background color.
-The shape is selected from a predefined SVG library based on its size and proportions, with minor adjustments for fitting.
+Taglines grouped by size and position use the same font and color.
+For the underlay, the color is chosen based on its contrast with the tagline color and similarity to the background color.
+The shape is selected from a predefined Scalable Vector Graphics (SVG) library based on its size and proportions, with minor adjustments for fitting.
 5. Experiments
@@
 We train the model using AdamW optimizer with a learning rate of 1e-5.
-Batch size on each GPU is set to 1 and gradient accumulation of 8 is adopted.
+Batch size on each GPU is set to 1, and gradient accumulation of 8 is adopted.
 The training costs about 14 hours with 8 NVIDIA A100 GPUs.
 During inference, we use top-p sampling with p set to 0.9 and a sampling temperature of 0.6.
-For background image generation, we use SDXL base model as the backbone.
+For background image generation, we use the SDXL base model as the backbone.
 During the first step, about 6.8 million high-aesthetic internal images with layout annotations are used.
 We train the newly added layers with a learning rate of 5e-5 for 20k steps on 512x512 resolution and further fine-tune them with a learning rate of 2e-5 for 20k steps on 1024x1024 resolution.
-Then in step 2, we introduce an inpainting ControlNet which is pre-trained with millions of advertising images from E-commerce platforms (Creative, 2024), and fine-tune the layout-related layers with a learning rate of 2e-5 for 1k steps, using 12k training data in the collected PIL dataset.
+Then in step 2, we introduce an inpainting ControlNet, which is pre-trained with millions of advertising images from E-commerce platforms (Creative, 2024), and fine-tune the layout-related layers with a learning rate of 2e-5 for 1k steps, using 12k training data in the collected PIL dataset.
 5.2. Evaluation Metrics
+To evaluate the overall poster generation and each module, we use the following metrics.
 Overall pipeline.
-We evaluate the effectiveness of the overall pipeline from two perspectives: overall visual quality and layout.
+We evaluate the overall pipeline from three perspectives: overall visual quality, and layout.
 For visual quality, we employ the Fréchet Inception Distance (FID) (Heusel et al., 2017) and aesthetic score.
-The aesthetic score is obtained by advertising experts ranking the outputs of different pipelines.
-The average rank is calculated to represent the mean aesthetic rating, where a lower rank indicates superior creative results.
+The aesthetic score comes from advertising experts ranking the outputs of different methods.
+The average rank represents the mean aesthetic rating, with a lower rank indicating better creative results.
 For layout, we adopt the metrics used in previous studies (Zhou et al., 2022; Seol et al., 2024), which are detailed in the subsequent Layout Generation module.
-Prompt generation. We evaluated prompt generation quality based on fore-background matching and e-commerce domain score.
-  • FBM Rate (Fore-background matching rate): It refers to how well a product foreground fits with a given prompt background.
-We classify this matching into “reasonable” or “unreasonable” through human annotation, and then determine the proportion of reasonable cases for each method.
-  • ED Score (E-commerce domain score): To assess how well the model-generated prompts reflect e-commerce characteristics, we calculate the FID between the CLIP features of these prompts and the ground truth prompts in the test set.
+Prompt generation.
+We evaluate prompt generation quality based on fore-background matching rate (FBM rate) and e-commerce domain score (ED score).
+FBM Rate refers to how well a product foreground fits with a given prompt background.
+We classify this matching into “reasonable” or “unreasonable” through human annotation, and calculate the proportion of reasonable cases.
+ED Score assesses how well the model-generated prompts reflect e-commerce characteristics.
+We calculate the FID between the CLIP features of these prompts and the ground truth prompts in the test set.
+Figure 5. Visualization of advertising images designed by different methods.
 Layout generation.
@@
 $Und_{l}$ and $Und_{s}$ checks if underlays enhance non-underlay elements.
-Content metrics assess the harmony of layout with the background image, including utility $Uti$, occlusion $Occ$, and unreadability $Rea$.
+Content metrics assess the harmony of layout with the image, including utility $Uti$, occlusion $Occ$, and unreadability $Rea$.
 Uti evaluates space usage for graphic elements.
 Occ is the average overlapping area between the graphic elements and products.
-Rea represents the non-flatness of regions that text elements without underlays are put on.
-For methods which take taglines as input, we report the text match rate (TMR), to see if the number of generated text bboxes are consistent with input taglines.
+Rea represents the non-flatness of regions that tagline elements without underlays are put on.
+For methods that take taglines as input, we report the tagline match rate (TMR), to see if the number of generated tagline bboxes is consistent with input taglines.
 Background image generation.
-We evaluate the quality of generated background images with FID, CLIP-T (Ruiz et al., 2023), CLIP-I (Gal et al., 2023).
-To verify layout control, we use Grounding-DINO (Liu et al., 2023b) to detect instances and calculate the maximum IoU between detected bboxes and ground-truth.
+We evaluate the quality of generated images with FID, CLIP-T (Ruiz et al., 2023), CLIP-I (Gal et al., 2022).
+To verify layout control, we use Grounding-DINO (Liu et al., 2024c) to detect instances and calculate the maximum IoU between detected bboxes and ground-truth bboxes.
 The average of these maximum IoUs is called mIOU.
@@
 5.3. Comparison
-Compare with automatic pipelines for product advertising image design.
-Since there are no public automatic methods for advertising image design conditioned on the product foreground, taglines and target size, we design two pipelines for comparison.
-Pipeline 1 use GPT-4o (OpenAI, 2024) for prompt and layout generation based on the input conditions.
-We fine-tune the SDXL-based inpainting model on our PIL dataset to obtain the image generation model.
+Compare with automatic pipelines for advertising image design.
+Since there are no public methods for advertising image design conditioned on the product foreground, taglines, and target size, we design two pipelines for comparison.
+Pipeline 1 use GPT-4o (OpenAI, 2024) for prompt and layout generation.
+We also finetune the SDXL-based inpainting model on our PIL dataset.
 The only difference between this method and ours is the absence of layout control and the corresponding training stage.
 During inference, only the product foreground is controlled by the layout output of GPT-4o.
-Pipeline 2 first generate a background prompt according to the foreground image using GPT-4o.
-The image generation model is the same as in Pipeline 1, but during inference, the foreground is placed according to general athetic rules (see details in the Appendix).
+Pipeline 2 first generate a background prompt using GPT-4o.
+The image generation model is the same as in Pipeline 1, but during inference, the foreground is placed according to general aesthetic rules.
 Next, the image and taglines are processed by the SOTA layout generation model PosterLlama to generate a graphic layout.
 Last, the results of these two pipelines are processed through the same graphics rendering module as our method to produce the final image.
-Table 1 shows the quantitative comparison between these pipelines and our PAID framework.
+Table 1 shows the quantitative comparison.
 PAID outperforms in most metrics, except for slightly higher occlusion compared to Pipeline 1, which places the product centrally with a small bounding box.
@@
 Pipeline 2 fixes the product location, limiting adaptability to product shapes, whereas PAID adjusts layouts based on product shape for better visuals.
-Additionally, the fixed background of Pipeline 2 can lead to layouts with crowded text boxes, making them hard to read, while our model provides more flexible and readable layouts.
+Additionally, the fixed background of Pipeline 2 can lead to layouts with crowded tagline boxes, making them hard to read, while our model provides more flexible and readable layouts.
 Table 1. Quantitative comparison between automatic pipelines for product-centric advertising image design.
@@
 PAID 37.524 1.405 0.0013 0.0017 0.9999 0.9973 0.1367 0.0955 0.1815
-Figure 7. Visualization of advertising image designed by different methods.
 Compare with prompt generation methods.
@@
 As shown in Table 2, our method outperforms the baselines in both FBM Rate and ED Score.
-We also use t-SNE for dimensionality reduction to visualize the feature distribution of prompts generated by our model before and after training (see Figure 8).
-This shows that our model effectively aligns with the e-commerce prompt dataset distribution after training.
-Table 2. Evaluation of prompt generation methods
+This means that our model effectively aligns with the e-commerce prompt dataset distribution after training.
+We also visualize the feature distribution of prompts generated in the appendix.
+Table 2. Evaluation of prompt generation methods.
 Method FBM Rate $\uparrow$ ED Score $\downarrow$
@@
 Finetuned XCP2 98.5% 7.17
-Figure 8. T-sne visualization of generated prompts.
 Compare with layout generation methods.
 To verify the advantage of our layout generation model, we compare it with previous background-based methods (DS-GAN (Hsu et al., 2023), RADM (Li et al., 2023b), and PosterLlama (Seol et al., 2024)) and the foreground-based method P&R (Li et al., 2023a).
-For background-based generation methods, we conduct experiments on two versions of the testing set for fair comparison, named the erased set and the generated set.
-The erased set contains images with graphic elements erased by LaMa (Suvorov et al., 2022), which may provide hints to put graphic elements on these erased area.
-The generated set contain images re-generated using their ground-truth prompts (image captions).
+For background-based generation methods, we conduct experiments on two versions of the testing set for fair comparison, named the erased/generated set.
+The erased set contains images with graphic elements erased by LaMa (Suvorov et al., 2022), which may provide hints to put graphic elements on these erased areas.
+The generated set contains images re-generated using their ground-truth prompts (image captions).
 The generated set has no above hints and conforms to real applications.
-As for the foreground-based method P&R, since it has not released the codes, we re-implement its core idea: first, generating the layout of the foreground and graphic elements conditioned on the foreground and taglines, and second, generating the final image based on the background prompt, layout, and product foreground.
-Our method uses the ground-truth prompt, product foreground, and taglines in PITA for prediction.
+As for the foreground-based method P&R, since it has not released the codes, we re-implement its core idea: first, generating the layout of the foreground and graphic elements conditioned on the foreground and taglines, and second, generating the final image based on the background prompt, layout, and foreground.
+Our method uses the ground-truth prompt, product foreground, and taglines for prediction.
 The results are shown in Table 3 and Figure 6.
@@
 This may be due to the fixed background image content, leaving less room for the model flexibility.
-RADM and DS-GAN, which are insensitive to tagline content, also face issues with generating text boxes that cannot adequately fill the tagline content.
+RADM and DS-GAN, which are insensitive to tagline content, also face issues with generating tagline boxes that cannot adequately fill the tagline content.
 The foreground-based method P&R improves in these areas but does not consider the harmony between the product, background prompt, and graphic layout, sometimes leading to unreasonable generated images.
@@
 Table 3.
-Quantitative performance comparison of layout generation methods on PITA dataset.
-Bold number represents the best result in each column except for erased set.
-BG-based and FG-based are the abbreviations for Background-based and Foreground-based methods, respectively.
+Quantitative performance comparison of layout generation methods.
+The bold number represents the best result in each column except for the erased set.
+BG/FG-based are the abbreviations for Background/Foreground-based methods.
 […]
@@
 Ours 1.0 0.0012 0.0017 0.9976 0.9956 0.1364 0.0973 0.1968 1.0
-Figure 9. Qualitative comparison of layout generation methods on PITA test set. Orange: text, Yellow: underlay, Red: logo, Black: Invalid
+Figure 6. Qualitative comparison of layout generation methods. Orange: tagline, Yellow: underlay, Red: logo, Black: Invalid.
 Compare with layout-controlled inpainting models.
@@
 Ours 25.917 0.313 0.906 0.705 0.079/0.127/0.078
-5.4. Ablation Studies
-Effect of JPGNL in layout generation.
+5.4. Online Result
+To assess the online performance of PAID, we conduct A/B tests in two advertising recommendation scenarios on Taobao.
+We randomly select 5,000 products from various categories (including clothing, cosmetics, food, and electronics) and collect their foreground images and taglines from the advertisers.
+We then generate advertising images for each product using both PAID and Pipeline 2 (the top-performing baseline as shown in Table 1).
+The A/B tests are conducted using 5% of the main traffic, affecting only the experimental products when this traffic is directed to them.
+After gathering data over one month, we found that PAID achieved a 3.02% and 3.03% increase in click-through rate (CTR) for the two recommendation scenarios, respectively.
+This indicates that PAID is capable of producing more visually appealing advertising images, resulting in improved recommendation outcomes.
+6. Conclusion
+In this paper, a novel framework named Product-Centric Advertising Image Design (PAID) is proposed to automatically create advertising images only based on product foreground images, marketing taglines, and target sizes.
+PAID consists of four stages: prompt generation, layout generation, background image generation, and graphics rendering.
+First, it generates adaptive prompts according to the product foregrounds.
+Then, it predicts how graphic and nongraphic elements should be placed according to the prompt, product foreground, and taglines, creating a harmonious overall layout.
+Next, a layout-controlled inpainting model is utilized for background image generation.
+A graphics rendering module is then applied to get the final images.
+We train separate experts to conduct these sub-tasks.
+Two datasets are created for the convenience of training and testing.
+The results of the test set and the online A/B tests have proven that the proposed PAID framework generates more visually pleasing and attractive advertising images.
+Acknowledgements
+This work is supported by National Natural Science Foundation of China (NSFC No.62076031 and No.62076036) and Alibaba Research Intern Program.
+Appendix A Dataset Examples
+In this section, we provide examples from the PITA and PIL datasets, illustrated in Figures 8 and 9, respectively.
+The PITA dataset includes advertising images with marketing taglines.
+In addition to depicting the layout of both graphic and nongraphic elements as shown in the figures, we also label the image caption (prompt), tagline content, and product foreground masks to train the prompt and layout generation models.
+Conversely, the PIL dataset contains product images without taglines.
+We annotate the layout of nongraphic elements and product foreground masks for training the image generation model.
+Figure 8. Examples from the PITA dataset, with tagline, underlay, and logo highlighted in orange, yellow, and red rectangles, respectively.
+Figure 9. Examples in PIL dataset.
+Appendix B Visualization of Prompt Distribution
+We use t-SNE for dimensionality reduction to visualize the feature distribution of prompts generated by our model before and after training (see Figure 10).
+This shows that our model effectively aligns with the e-commerce prompt dataset distribution after training.
+Appendix C Ablation Studies
+To verify the effectiveness of our design, we conduct ablation studies on the methods and strategies used in each module.
+Figure 10. T-sne visualization of generated prompts.
+C.1. Effect of JPGNL in Layout Generation
 Our method predicts the overall layout of graphic and nongraphic elements.
 Here, we examine the need to predict the layout of nongraphic elements according to the background prompt.
-As shown in Figure 11, without specifying the position of nongraphic elements, texts placed on complex areas can lead to a messy appearance.
-By predicting the overall layout, we improve the text readability of advertising images.
+As shown in Figure 11, without specifying the position of nongraphic elements, taglines placed on complex areas can lead to a messy appearance.
+By predicting the overall layout, we improve the tagline readability of advertising images.
 Additionally, with nongraphic layout predictions, generated images align more closely with background prompts.
@@
 Besides, in Table 5, we quantitatively analyze the impact of predicting nongraphic layouts on graphic layouts.
-Predicting nongraphic layouts (w/ JPGNL) can improve metrics related to product occlusion and text readability.
+Predicting nongraphic layouts (w/ JPGNL) can improve metrics related to product occlusion and tagline readability.
 However, since it adds complexity to the task, it negatively affects graphic metrics such as element overlap and alignment.
-Figure 11. An illustration of the effect of JPGNL. Predicting image layout makes better visual effect.
 Table 5. Quantitative ablation study on JPGNL.
@@
 w/ JPGNL 1.0 0.0012 0.0017 0.9976 0.9956 0.1364 0.0973 0.1968
-Effect of RKBR and CCLP in layout generation.
+C.2. Effect of RKBR and CCLP in Layout Generation
 We validate the effectiveness of the proposed input and output format for layout generation, including RKBR and CCLP.
@@
 Adding CCLP improves the Uti and Occ values on No Occ Set, indicating that CCLP helps the model to distinguish occlusion-allowing sets and others by providing explicit class conditions.
+Figure 11. An illustration of the effect of JPGNL. Predicting image layout makes better visual effect.
 Table 6.
@@
 1.0 0.0012 0.0017 0.9976 0.9956 0.1364/0.1493/0.0849 0.0973/0.0136/0.4316 0.1968 1.0
-Effect of DLC and LAT in layout-to-image model.
+C.3. Effect of DLC and LAT in Layout-to-Image Model
 We investigate the effect of deep layer control (DLC) and Lora adaptation training (LAT), with results shown in Table 7.
@@
 25.917 0.313 0.906 0.705 0.079/0.127/0.078
-6. Conclusion
-In this paper, a novel framework named Product-Centric Advertising Image Design (PAID) is proposed to automatically create ad images only based on product foreground images, marketing taglines, and target sizes.
-PAID consists of four serial stages: prompt generation, layout generation, background image generation, and graphics rendering.
-First, it generates adaptive prompts according to the semantic and visual information of the product, for the benefit of background image generation.
-Then, it predicts how graphic and nongraphic elements should be placed according to the prompt, product foreground, and taglines, creating a harmonious overall layout.
-Next, a layout-controlled inpainting model is utilized for background image generation.
-A graphics rendering module is then applied to get the final images.
-We train separate experts to conduct these sub-tasks.
-Two high-aesthetic and diverse product-centric datasets are created for the convenience of training.
-Qualitative and quantitative comparisons prove that the proposed PAID framework generates more visually pleasing ad images, compared with previous methods.
-Acknowledgements
-This work is supported by National Natural Science Foundation of China (NSFC No.62076031 and No.62076036) and Alibaba Research Intern Program.

### v3

@@
-PAID: A Framework of Product-Centric Advertising Image Design
+T-Stars-Poster: A Framework for Product-Centric Advertising Image Design
 Abstract.
-Creating visually appealing advertising images is often a labor-intensive and time-consuming process.
-Is it possible to automatically generate such images using only basic product information—specifically, a product foreground image, taglines, and a target size?
-Existing methods mainly focus on parts of the problem and fail to provide a comprehensive solution.
-To address this gap, we propose a novel multistage framework called Product-Centric Advertising Image Design (PAID).
+Creating advertising images is often a labor-intensive and time-consuming process.
+Can we automatically generate such images using basic product information like a product foreground image, taglines, and a target size?
+Existing methods mainly focus on parts of the problem and lack a comprehensive solution.
+To bridge this gap, we propose a novel product-centric framework for advertising image design called T-Stars-Poster.
 It consists of four sequential stages to highlight product foregrounds and taglines while achieving overall image aesthetics: prompt generation, layout generation, background image generation, and graphics rendering.
-Different expert models are designed and trained for the first three stages: First, we use a visual language model (VLM) to generate background prompts that match the products.
+Different expert models are designed and trained for the first three stages: First, a visual language model (VLM) generates background prompts that match the products.
 Next, a VLM-based layout generation model arranges the placement of product foregrounds, graphic elements (taglines and decorative underlays), and various nongraphic elements (objects from the background prompt).
-Following this, we train an SDXL-based image generation model that can simultaneously accept prompts, layouts, and foreground controls.
-To support the PAID framework, we create corresponding datasets with over 50,000 labeled images.
-Extensive experimental results and online A/B tests demonstrate that PAID can produce more visually appealing advertising images.
+Following this, an SDXL-based model can simultaneously accept prompts, layouts, and foreground controls to generate images.
+To support T-Stars-Poster, we create two corresponding datasets [^1] The primary dataset is available at https://github.com/alimama-creative/PITA-dataset.
+with over 50,000 labeled images.
+Extensive experiments and online A/B tests demonstrate that T-Stars-Poster can produce more visually appealing advertising images.
 Figure 1. Generated advertising images by our methods with a product foreground image, taglines, and a target size as input.
@@
 Advertising images (as shown in Figure 1) are essential for commercial recommendation.
-Visually appealing and attractive images usually have high click-through rates (Wang et al., 2021) but are labor-intensive and time-consuming to create (Vashishtha et al., 2024).
+Visually appealing images usually have high click-through rates (Wang et al., 2021) but are labor-intensive and time-consuming to create (Vashishtha et al., 2024).
 With the rapid advancement of generative methods (Menick and Kalchbrenner, 2018; Van Den Oord et al., 2016; Ho et al., 2020; Rombach et al., 2021), it is now possible to automatically generate advertising images using only basic product inputs: a product foreground image, marketing taglines, and a target size.
-Currently, generating advertising images using only the basic inputs mentioned above has rarely been explored.
+Currently, generating advertising images using only the basic inputs has rarely been explored.
 Existing methods (Du et al., 2024; Weng et al., 2024; Wang et al., 2024b; Lin et al., 2023; Zhou et al., 2022) mainly focus on parts of the problem, such as background or layout generation, rather than offering a complete solution for integrating product foregrounds and taglines.
 Although these methods can be combined, they struggle to effectively highlight product images and taglines while maintaining visual appeal and harmony.
-Specifically, building on existing approaches (Du et al., 2024; Weng et al., 2024; Wang et al., 2024b; Lin et al., 2023; Zhou et al., 2022), we can first utilize predefined rules to create background prompts.
-These prompts are then fed into text-to-image inpainting models to complete backgrounds for product foregrounds.
+Specifically, building on existing approaches (Du et al., 2024; Weng et al., 2024; Wang et al., 2024b; Lin et al., 2023; Zhou et al., 2022), we can first utilize predefined rules to create background prompts, which are then fed into text-to-image inpainting models to complete backgrounds for product foregrounds.
 Layouts for taglines and decorative elements (collectively referred to as graphic elements) are predicted based on the images and taglines.
 Finally, attribute prediction and graphics rendering techniques are employed to add these graphic elements to the images.
-However, this process has several limitations: a) Lack of adaptability: It cannot generate background prompts that match the specific content and shape of the product, leading to poor foreground-background harmony.
-b) Fixed foreground positioning: The position of the foreground cannot be adjusted based on its characteristics and the background prompt, potentially resulting in inappropriate product size and poor composition.
-c) Restricted tagline placement: When adding taglines, the whole image content is already set, limiting the available space for tagline placement.
-Figure 2. Pipeline of PAID. It consists of four stages and generates advertising images centered around product information.
-In this paper, we propose a novel framework called Product-Centric Advertising Image Design (PAID) to enhance advertising image generation using only basic product inputs.
-While PAID remains a multistage process, we have refined its stages and task definitions to enhance foreground-background compatibility and the spatial arrangement of product foregrounds and taglines.
-As illustrated in Figure 2, we first introduce a task to generate prompts, using a Visual Language Model (VLM) (Dong et al., 2024) to adaptively create background prompts that better match the scene, placement, and angle based on product foregrounds.
-Following this, we propose a layout generation task that arranges the graphic elements (marketing taglines and decorative underlays), product foreground, and other nongraphic elements (objects to be generated from the background prompt).
-This arrangement is based on the product foreground, background prompt, marketing taglines, and target size.
-By arranging the product foreground and marketing information concurrently, as opposed to generating the overall image first and then deciding the graphic layout, it becomes less constrained by the image content and allows for a more rational presentation of the product info.
-Furthermore, considering the nongraphic elements together helps minimize conflicts with taglines that could affect readability and attractiveness.
-Next, we train an inpainting text-to-image model based on SDXL (Podell et al., 2023) that can be controlled by input layouts.
+However, this process has several limitations: a) Lack of adaptability: The background prompts cannot match the specific content and shape of the product, leading to poor foreground-background harmony.
+b) Fixed foreground positioning: The foreground position cannot be adjusted based on the product characteristics and the background prompt, potentially resulting in inappropriate product size and poor composition.
+c) Restricted tagline placement: The whole image content is set before taglines are added, limiting space for tagline placement.
+Figure 2. Pipeline of T-Stars-Poster. It consists of four stages and generates advertising images centered around product information.
+In this paper, we propose a novel product-centric framework called T-Stars-Poster for advertising image generation using only basic product inputs.
+We refine the stages and task definitions to enhance foreground-background compatibility and the spatial arrangement of product foregrounds and taglines.
+As illustrated in Figure 2, we first introduce a task to generate prompts, using a Visual Language Model (VLM) (Dong et al., 2024) to create background prompts that match the scene, placement, and angle with product foregrounds.
+Next, a layout generation task arranges the graphic elements (marketing taglines and decorative underlays), product foreground, and other nongraphic elements (objects to be generated from the background prompt).
+This arrangement is based on the product foreground, background prompt, taglines, and target size.
+By arranging the product foreground and taglines simultaneously, rather than generating the whole image before deciding the graphic layout, we reduce layout constraints and present product information more effectively.
+Furthermore, considering nongraphic elements together helps minimize conflicts with taglines that could affect readability and attractiveness.
+We then train an inpainting text-to-image model based on SDXL (Podell et al., 2023) that can be guided by input layouts.
 Combining inpainting, layout, and prompt control effectively is challenging due to potential conflicts among multiple control signals.
 A LoRA (Hu et al., 2021) adaption strategy is adopted to balance conflicts and facilitate training.
-Since layout mainly influences higher-level abstract information, we apply layout control only in the deeper blocks of the UNet (He et al., 2016), combined with a multi-scale training strategy to enhance training stability and efficiency.
-Finally, we utilize an attribute prediction and graphics rendering module to overlay graphic elements onto the image.
-To build the framework and verify its effectiveness, we have collected and labeled over 50,000 images to establish datasets.
-In testing, PAID demonstrates its ability to produce advertising images with better visual quality.
+Since layout mainly influences higher-level abstract information, we apply layout control only in the deeper blocks of the UNet (He et al., 2016), along with a multi-scale training strategy to enhance stability and efficiency.
+Finally, we use an attribute prediction and graphics rendering module to overlay graphic elements onto the image.
+To build the framework and verify its effectiveness, we have collected and labeled over 50,000 images for datasets.
+In test, T-Stars-Poster shows it can produce advertising images with better visual quality.
 It has been deployed on two advertising recommendation scenarios of Taobao, and online A/B tests show that its generated images are more appealing to users.
 In summary, our main contributions are as follows:
-  • We propose a novel framework PAID for advertising image design given only a product foreground image, taglines, and a target size.
-It centers on the product foreground and taglines to enhance the overall product prominence and improve the image aesthetics.
+  • We propose a framework T-Stars-Poster for advertising image design given only a product foreground, taglines, and a target size.
+It centers on product foregrounds and taglines to improve the product prominence and image aesthetics.
   • A VLM is introduced to generate suitable prompts for image generation using the input product foreground image, considering product placement angle and shape.
@@
 2.1. Advertising Image Design
-In recent years, many methods have been proposed to automate advertising image design (Zhou et al., 2022; Hsu et al., 2023; Li et al., 2023b; Seol et al., 2024; Lin et al., 2023; Li et al., 2023a).
-They can be divided into two subcategories: background-based and foreground-based methods.
-Background-based methods predict graphic layouts based on complete images (not only product foregrounds), followed by attribute prediction and graphics rendering modules to create the final advertising image.
-CGL-GAN (Zhou et al., 2022) and DS-GAN (Hsu et al., 2023) design transformer architectures to leverage image info and generate layouts.
-RADM (Li et al., 2023b) is the first to simultaneously consider the image and tagline content to derive layouts with a diffusion model.
-PosterLlama (Seol et al., 2024) leverages the rich design knowledge in Large Language Models (LLMs) to predict the layout.
-AutoPoster (Lin et al., 2023) introduces a complete pipeline which first generates a layout based on a complete image, and then predicts the attributes of graphic elements and renders them onto the image.
-For these methods, the spatial layout is restricted due to fixed image content, which further restricts the tagline content, including the number and length of taglines.
-In contrast, P&R (Li et al., 2023a) is a foreground-based method that leverages both the product foreground image and tagline content to predict the layout of the product foreground and taglines.
-It subsequently designs a module network to outpaint the background and render the taglines according to the given background prompt and generated layout.
-The tagline content is not limited, and the layout is more flexible.
+Many methods have been proposed to automate advertising image design (Zhou et al., 2022; Hsu et al., 2023; Li et al., 2023b; Seol et al., 2024; Lin et al., 2023; Li et al., 2023a), falling into two categories: background-based and foreground-based.
+Background-based methods (Zhou et al., 2022; Hsu et al., 2023; Li et al., 2023b; Seol et al., 2024; Lin et al., 2023) predict graphic layouts for complete images (not product foregrounds), followed by attribute prediction and graphics rendering to create the final image.
+These methods face spatial constraints due to fixed image content, limiting tagline content.
+In contrast, P&R (Li et al., 2023a) is a foreground-based method using both foreground images and tagline content to predict layout, then outpainting the background and rendering taglines according to the given background prompt and generated layout.
+The tagline content and layout are more flexible.
 However, P&R ignores the harmony between the background prompt and the generated layout.
 The layout may conflict with background generation.
-In our PAID framework, the overall layout of graphic elements, product foreground, and other nongraphic elements is generated jointly according to the product foreground, background prompt, and tagline contents, aiming to achieve the best harmony among them.
+Our framework generates the overall layout for graphic elements, product foreground, and other elements jointly, using product foregrounds, background prompts, and tagline content to ensure harmony.
 2.2. Controllable Image Generation
 There have been significant developments in text-to-image models (Rombach et al., 2021; Podell et al., 2023).
-To control instances and further improve image quality, some methods introduce layout guidance, dividing into training-free and training-based categories.
-Training-free methods manipulate the attention map (Kim et al., 2023) or compute attention-based loss during de-noising (Chefer et al., 2023; Xie et al., 2023; Chen et al., 2024; Bar-Tal et al., 2023).
-Training-based methods, like GLIGEN (Li et al., 2023c) and InstanceDiffusion (Wang et al., 2024a), modify network architecture, incorporating layout data by adding gated self-attention layers.
-For advertising image generation (Du et al., 2024; Wang et al., 2023; Cao et al., 2024; Li et al., 2023a), text-to-image models equipped with inpainting capabilities are frequently utilized.
-The inpainting technique is employed to preserve product characteristics, while the prompt is used to describe the background.
-The foreground image is utilized by incorporating extra channels into the UNet architecture (Rombach et al., 2021) or employing an inpainting ControlNet (Zhang et al., 2023).
-To better control the overall layout of advertising images and improve the quality of generated images, we combine inpainting with layout control.
+To control instances and improve image quality, some methods introduce layout guidance, dividing into training-free and training-based categories.
+Training-free methods adjust the attention map or compute attention-based loss during inference (Kim et al., 2023; Chefer et al., 2023; Xie et al., 2023; Chen et al., 2024; Bar-Tal et al., 2023).
+Training-based methods (Li et al., 2023c; Wang et al., 2024a) add gated self-attention layers to models and train them with corresponding data.
+For advertising image generation (Du et al., 2024; Wang et al., 2023; Cao et al., 2024; Li et al., 2023a), text-to-image models with inpainting capabilities are often utilized.
+Inpainting preserves foregrounds and prompts describe the backgrounds.
+The foreground image is integrated using extra channels (Rombach et al., 2021) or an inpainting ControlNet (Zhang et al., 2023).
+To better manage the layout and enhance image quality, we combine inpainting with layout control.
 Currently, there is limited exploration in this area.
-SceneBooth (Chai et al., 2025) is the first model to integrate both controls in training, combining gated self-attention layers and inpainting ControlNet to train a model based on SD 1.5 (Rombach et al., 2021) that can accept layout and foreground controls.
+SceneBooth (Chai et al., 2025) is the first to integrate both controls in training, combining gated self-attention layers and inpainting ControlNet to train a model based on SD 1.5 (Rombach et al., 2021).
 However, directly combining the two to train a larger model like SDXL can lead to non-convergence or a decline in the quality of generated images.
-Therefore, we use several strategies when training the layout-controlled inpainting model to ensure training stability, generation quality, and efficiency.
+Therefore, we use several strategies to ensure training stability, generation quality, and efficiency.
 3. PITA Dataset and PIL Dataset
-For automatic image design of product-centric advertising, we collect a large-scale Product-Centric Image-Tagline Advertising (PITA) dataset, gathering 38,017 samples from e-commerce platforms and the CGL dataset (Zhou et al., 2022).
-1,000 of them are used for testing.
-PITA is a dataset of varying sizes, featuring images with four distinct aspect ratios: 0.684, 1.0, 0.667, and 0.75.
-Each image is labeled with prompts (foreground caption and background caption), a product mask, and a layout of graphic and nongraphic elements.
+We collect a Product-Centric Image-Tagline Advertising (PITA) dataset with 38,017 samples from Taobao and the CGL dataset (Zhou et al., 2022).
+1,000 of them are for test.
+PITA covers major categories of Taobao, such as clothing, food, furniture, and electronics.
+It features images with four distinct aspect ratios: 0.684, 1.0, 0.667, and 0.75.
+In the collection process, we exclude images with messy backgrounds, plain colors without shadows, unattractive stickers, close-ups, poorly presented taglines, or cluttered elements.
+Each has labels for prompts (foreground and background captions), a product mask, and a layout.
 Each element is represented with a type and a bounding box (bbox).
-The graphic elements contain “Logo”, “Tagline”, and “underlay”.
-Some examples of data visualization are in the appendix.
-For layout-controlled background inpainting, we additionally construct a Product-Centric Image Layout (PIL) dataset.
-The images do not contain graphic elements and are labeled with prompts, product masks, and layouts of nongraphic elements.
-PIL contains 12,247 samples in total and 1,000 of them are divided for test.
+Graphic elements contain “Logo”, “Tagline”, and “Underlay”.
+Initial annotations are generated using automatic methods such as DAMO Academy’s matting API (Academy, 2024) for foreground extraction, GPT-4o (OpenAI, 2024) for image captioning, a detection model (Dai et al., 2021) trained with the CGL dataset and Grounding-DINO (Liu et al., 2024c) for element detection, and an OCR model (PaddlePaddle, 2023) for tagline extraction.
+Annotations are then manually reviewed and corrected for accuracy.
+We also develop a Product-Centric Image Layout (PIL) dataset for layout-controlled background inpainting.
+It contains images without graphic elements and includes 12,247 samples, with 1,000 reserved for testing.
+The filtering and annotation process is the same as the PITA dataset, excluding graphic elements.
 4. Method
 4.1. Overall Framework
-As illustrated in Figure 2, our proposed PAID framework consists of four stages: prompt generation, layout generation, background image generation, and graphics rendering.
+As illustrated in Figure 2, our proposed T-Stars-Poster framework consists of four stages: prompt generation, layout generation, background image generation, and graphics rendering.
 4.2. Prompt Generation
 For prompt generation, we fine-tune a large vision-language model $\pi_{prompt}$ based on InternLM-XComposer2-vl (XCP2) (Dong et al., 2024). The method is illustrated in the upper part of Figure 3.
-Input data format.
-For the input of LLM, we first describe the prompt generation task and ask for foreground and background descriptions.
-For the image input, we crop out the product foreground and feed it into the visual encoder.
+Input data format. We describe the prompt generation task and request foreground/background descriptions. Meanwhile, we feed the product foreground into the visual encoder.
 Output data format.
-To enable downstream layout and background image generation tasks, and to better understand the relationship between the foreground and background, we use the prompt generation model $\pi$ to simultaneously predict the foreground description $p_{fore}$ and the background prediction $p_{back}$.
+To support layout and background image generation tasks and understand the foreground-background relationship, the prompt generation model $\pi$ predicts both the foreground description $p_{fore}$ and background description $p_{back}$ simultaneously.
 We format the output in JSON as it is compatible with the pre-trained VLM and simplifies the following analysis.
@@
 Training scheme.
-A standard VLM training scheme involves two stages: pre-training and supervised fine-tuning (SFT) (Liu et al., 2024a).
-During pre-training, the projector learns to align features between visual and text modalities.
-In the SFT stage, the projector and LLM are further trained for visual understanding tasks.
-Since we utilize the VLM trained after the SFT stage, the model can understand input images and follow instructions.
-We only fine-tune the LLM branch with our dataset to adapt the model to the prompt generation task, using cross-entropy loss as the objective function.
-The model $\pi_{prompt}$ takes the foreground image $I_{fore}$ and predefined instructions as input, performing two main tasks: generating a foreground description $p_{fore}$ and predicting the background description $p_{back}$.
+We use a VLM trained after supervised fine-tuning (SFT) (Liu et al., 2024a), allowing the model to understand images and follow instructions.
+We only fine-tune the LLM part with our dataset for the prompt generation task, using cross-entropy loss.
+The model $\pi_{prompt}$ takes a foreground image $I_{fore}$ and predefined instructions to generate a foreground description and predict a background description.
 The prompt generation process is as follows:
-(1) $p=\pi_{prompt}(I_{fore}).$
+(1) $p=(p_{fore},p_{back})=\pi_{prompt}(I_{fore}).$
 4.3. Layout Generation
@@
 The method is shown in the lower part of Figure 3.
-Figure 3. The framework of our prompt and layout generation model and data format.
+Figure 3. The framework of prompt and layout generation.
 Input data format.
-For the input to the LLM, we describe the layout generation task and organize input conditions in JSON format.
-The image input remains the same as in prompt generation.
-A target canvas aspect ratio is given as a condition to support multi-scale layout design.
-The aspect ratio of the product foreground is provided to help keep its shape.
-Foreground and background prompts are also given for generating the image layout.
-For the pre-defined taglines, we list them with empty locations and expect the model to complete them.
-If a logo is needed, we include it as a condition in the graphic layout section and specify its aspect ratio.
-Except for these basic conditions, we also take product characteristics into account.
-In E-commerce scenarios, some products should not be occluded for better display.
-While for others, like clothing or human-hold items, partial occlusion is acceptable.
-To differentiate between these, we introduce a Class-Conditioned Layout Prediction (CCLP) strategy.
-CCLP provides clear information about the product class and whether foreground occlusion is allowed.
+For LLM input, we describe the layout generation task and organize conditions in JSON format.
+The image input is the same as in prompt generation.
+A target canvas aspect ratio is provided to support multi-scale design.
+The aspect ratio of the product foreground is given to keep its shape.
+Foreground/background prompts are included for image layout generation, and taglines are listed with empty locations for the model to fill in.
+If a logo is needed, it will be included with its aspect ratio.
+We also take product characteristics into account.
+In e-commerce, some products need to be fully visible, while partial occlusion is acceptable for others, like clothing.
+To address this, we use a Class-Conditioned Layout Prediction (CCLP) strategy that indicates the product class and whether occlusion is allowed.
 For the not-allowing occlusion set (No Occ Set), we describe “The class of subject is [V1].
@@
 The output layout is structured in JSON format with three parts: subject layout, other nongraphic layout, and graphic layout.
-To maintain the shape of the product foreground, we propose a Ratio-Keeping Bbox Representation (RKBR).
-It is expressed as [x, y, h*r1/r2, h].
+To maintain the foreground shape, we propose a Ratio-Keeping Bbox Representation (RKBR): [x, y, h*r1/r2, h].
 (x,y) is the normalized center point.
@@
 r1 and r2 are the actual aspect ratio of the foreground and canvas, respectively.
-Each value ranges from 0 to 1 with three decimal places.
+Each value ranges from 0 to 1.
 The nongraphic layout contains the location of elements in the background prompt, each having a type (instance name) and bounding box (bbox).
-The graphic layout defines each element by type, content, and bounding box, covering tagline, underlay, and logo.
+The graphic layout defines each element by type, content, and bounding box, covering the tagline, underlay, and logo.
 The content key is used only for tagline types to specify its semantic content.
@@
 The model is expected to supplement underlay adaptively.
-If logos are required, it is represented in the same way as the product for ratio keeping.
+If logos are required, it is represented in the same way as the foreground for ratio keeping.
 Training scheme.
-Same as the prompt generation module, we only fine-tune the LLM branch with the visual encoder and projector frozen.
-The output JSON is structured to predict the layout in a raster scan order, adhering to the ascending order of top-left coordinates.
-Furthermore, we shuffle the taglines in the input conditional JSON for data augmentation and prevent the model from memorizing the input sequence.
+Like the prompt generation module, we fine-tune only the LLM branch.
+The output JSON predicts layouts in a raster scan order, following ascending top-left coordinates.
+We also shuffle taglines in the input JSON for data augmentation to prevent the model from memorizing the sequence.
 4.4. Background Image Generation
@@
 Layout control.
-Suppose that there are $N$ elements to be drawn in an image, represented as $\{(n_{i},l_{i})|i=0,1,...,N-1\}$.
+Suppose an image has $N$ elements, represented as $\{(n_{i},l_{i})|i=0,1,...,N-1\}$.
 We map these into $N$ embeddings as layout input.
 Here, $n_{i}$ is the element name, and $l_{i}$ is its location.
-Concretely, we use the text encoder in SDXL to map $n_{i}$ into a global semantic embedding $\textbf{e}_{n,i}$.
+We use the text encoder in SDXL to map $n_{i}$ into a global semantic embedding $\textbf{e}_{n,i}$.
 $l_{i}$ is transferred into normalized top-left and bottom-right coordinates, then fed into a Fourier mapping layer(Tancik et al., 2020), deriving the location embedding $\textbf{e}_{l,i}$.
@@
 A gated self-attention (GSA) layer (Li et al., 2023c) is inserted between the self- and cross-attention layers in the original UNet to inject layout embeddings.
-Supposing there are $M$ visual embeddings (V) after the self-attention layer.
-The output of the GSA layer is
+If there are $M$ visual embeddings (V) after the self-attention layer, the GSA output is
 $GSA(\textbf{V},\textbf{E})=\textbf{V}+tanh(\gamma)*SA(cat(\textbf{V},\textbf{E}))[:M]$,
-where $\gamma$ is a learnable parameter, $cat$ means concatenation, and $SA$ means self-attention.
-Considering that the layout is semantic-level information and less related to high-frequency details, we only add GSA layers in the deep blocks of UNet.
+where $\gamma$ is a learnable parameter, catcat means concatenation, and $SA$ means self-attention.
+Since layout relates to semantic-level information and not high-frequency details, GSA layers are only added in the deep blocks of UNet.
 We refer to this strategy as Deep Layer Control (DLC).
 Specifically, the layout control is applied in the middle block and the lowest resolution of up blocks.
-This reduces the parameters and inference latency, and has no negative impact on generation quality.
-Additionally, we adopt a LoRA Adaptation Training (LAT) strategy in which trainable LoRA (Hu et al., 2021) layers are added to UNet, to bridge the gap between the base model and added GSA layers.
-This further accelerates convergence and increases generation quality.
+Additionally, we use a LoRA Adaptation Training (LAT) strategy, adding LoRA (Hu et al., 2021) layers to UNet to bridge the gap between the base model and added GSA layers.
+These further accelerate convergence and increase generation quality.
 Figure 4 describes how layout control is achieved.
@@
 4.5. Graphics Rendering
-Since there is no effective method to predict the visual attributes of graphic elements, and the graphics rendering module is not the focus of this paper, we design a strategy according to the aesthetic rules to handle graphics rendering.
-We first sort the tagline elements by area to decide the font and color, both of which are limited to a specific set (Liu et al., 2024b).
-The selection is based on the similarity to the foreground color and contrast with the background area.
+Since there is no effective method to predict the visual attributes of graphic elements, and the graphics rendering module is not the focus of this paper, we design a rendering strategy.
+We sort taglines by area to select the font and color from a limited set (Liu et al., 2024b).
+The choice depends on the similarity to the foreground color and contrast with the background.
 Taglines grouped by size and position use the same font and color.
-For the underlay, the color is chosen based on its contrast with the tagline color and similarity to the background color.
-The shape is selected from a predefined Scalable Vector Graphics (SVG) library based on its size and proportions, with minor adjustments for fitting.
+The underlay color is picked for contrast with the tagline and similarity to the background, while the shape is selected from a predefined SVG library based on size and proportions, with slight modifications to fit.
 5. Experiments
-In this section, we compare our PAID framework and its modules with previous ones. We conducted experiments on the PITA dataset, and the results confirm the effectiveness of our method.
+In this section, we compare our framework with previous models.
 5.1. Implementation Details
-For prompt and layout generation, XCP2 (Dong et al., 2024) is used as our backbone.
-Its visual encoder is built upon CLIP ViT-Large (Radford et al., 2021), processing 490x490 resolution images.
-The LLM is based on InternLM2 (Cai et al., 2024).
-We fine-tune the LLM branch on the PITA dataset for one epoch.
-We train the model using AdamW optimizer with a learning rate of 1e-5.
-Batch size on each GPU is set to 1, and gradient accumulation of 8 is adopted.
-The training costs about 14 hours with 8 NVIDIA A100 GPUs.
-During inference, we use top-p sampling with p set to 0.9 and a sampling temperature of 0.6.
-For background image generation, we use the SDXL base model as the backbone.
-During the first step, about 6.8 million high-aesthetic internal images with layout annotations are used.
-We train the newly added layers with a learning rate of 5e-5 for 20k steps on 512x512 resolution and further fine-tune them with a learning rate of 2e-5 for 20k steps on 1024x1024 resolution.
-Then in step 2, we introduce an inpainting ControlNet, which is pre-trained with millions of advertising images from E-commerce platforms (Creative, 2024), and fine-tune the layout-related layers with a learning rate of 2e-5 for 1k steps, using 12k training data in the collected PIL dataset.
+We use XCP2 (Dong et al., 2024) as the backbone for prompt and layout generation, fine-tuning the LLM branch on the PITA dataset for one epoch respectively.
+The model is trained with the AdamW optimizer at a 1e-5 learning rate, using a total batch size of 64.
+Training takes 14 hours on 16 NVIDIA H20 GPUs.
+For inference, we apply top-p sampling with p set to 0.9 and a sampling temperature of 0.6.
+For background image generation, we use the SDXL model as the backbone.
+In step 1, we train new layers using 6.8 million aesthetic images (from the LAION (Schuhmann et al., 2022) and PIL dataset) with layout annotations at a learning rate of 5e-5 for 20k steps on 512x512 resolution, taking 65 hours on 16 NVIDIA H20 GPUs (batch size 64).
+We then fine-tune them with a learning rate of 2e-5 for 20k steps at 1024x1024 resolution over 86 hours (batch size 32).
+In step 2, we introduce an inpainting ControlNet pre-trained on advertising images (Creative, 2024).
+We fine-tune the layout-related layers with a learning rate of 2e-5 for 1k steps using 12k training data from the PIL dataset, which takes 5 hours with 16 NVIDIA H20 GPUs (batch size 16).
 5.2. Evaluation Metrics
-To evaluate the overall poster generation and each module, we use the following metrics.
+We use the following metrics to assess the pipeline and each module.
 Overall pipeline.
-We evaluate the overall pipeline from three perspectives: overall visual quality, and layout.
-For visual quality, we employ the Fréchet Inception Distance (FID) (Heusel et al., 2017) and aesthetic score.
-The aesthetic score comes from advertising experts ranking the outputs of different methods.
-The average rank represents the mean aesthetic rating, with a lower rank indicating better creative results.
-For layout, we adopt the metrics used in previous studies (Zhou et al., 2022; Seol et al., 2024), which are detailed in the subsequent Layout Generation module.
+We assess the pipeline from two perspectives: visual quality and layout.
+For visual quality, we use the Fréchet Inception Distance (FID) (Heusel et al., 2017) and an aesthetic score, with the latter based on rankings by advertising experts.
+A lower average rank indicates better results.
+For layout evaluation, we use metrics from previous work (Zhou et al., 2022; Seol et al., 2024), as detailed in the layout generation module.
 Prompt generation.
 We evaluate prompt generation quality based on fore-background matching rate (FBM rate) and e-commerce domain score (ED score).
-FBM Rate refers to how well a product foreground fits with a given prompt background.
+FBM Rate measures how well a product foreground fits with the background prompt.
 We classify this matching into “reasonable” or “unreasonable” through human annotation, and calculate the proportion of reasonable cases.
@@
 Layout generation.
-Following previous work (Zhou et al., 2022; Seol et al., 2024), we use graphic metrics and content metrics to evaluate layouts.
+Following previous work (Zhou et al., 2022; Seol et al., 2024), we use graphic/content metrics to evaluate layouts.
 Graphic metrics focus on relationships between graphic elements, including validity $Val$, alignment $Ali$, overlap $Ove$, and underlay $Und_{l},Und_{s}$.
-Val means the ratio of elements larger than 0.1% of the canvas.
-Ali assesses how well elements are spatially aligned.
-Ove is the average intersection over union (IoU) of all element pairs except for underlays.
-$Und_{l}$ and $Und_{s}$ checks if underlays enhance non-underlay elements.
 Content metrics assess the harmony of layout with the image, including utility $Uti$, occlusion $Occ$, and unreadability $Rea$.
-Uti evaluates space usage for graphic elements.
-Occ is the average overlapping area between the graphic elements and products.
-Rea represents the non-flatness of regions that tagline elements without underlays are put on.
-For methods that take taglines as input, we report the tagline match rate (TMR), to see if the number of generated tagline bboxes is consistent with input taglines.
+For methods that take taglines as input, we report the tagline match rate (TMR), to see if the number of generated tagline bboxes matches the input taglines.
 Background image generation.
 We evaluate the quality of generated images with FID, CLIP-T (Ruiz et al., 2023), CLIP-I (Gal et al., 2022).
-To verify layout control, we use Grounding-DINO (Liu et al., 2024c) to detect instances and calculate the maximum IoU between detected bboxes and ground-truth bboxes.
-The average of these maximum IoUs is called mIOU.
+To verify layout control, we use Grounding-DINO (Liu et al., 2024c) to detect instances and calculate the mIOU between detected bboxes and ground-truth bboxes.
 In addition, Grounding-DINO AP scores are also reported.
@@
 Compare with automatic pipelines for advertising image design.
-Since there are no public methods for advertising image design conditioned on the product foreground, taglines, and target size, we design two pipelines for comparison.
+Since there are no public methods for advertising image design conditioned on the product foregrounds, taglines, and target sizes, we design two pipelines for comparison.
 Pipeline 1 use GPT-4o (OpenAI, 2024) for prompt and layout generation.
-We also finetune the SDXL-based inpainting model on our PIL dataset.
-The only difference between this method and ours is the absence of layout control and the corresponding training stage.
+It fine-tunes an SDXL-based inpainting model on our PIL dataset.
+The main difference from our approach is the lack of layout control.
 During inference, only the product foreground is controlled by the layout output of GPT-4o.
@@
 Table 1 shows the quantitative comparison.
-PAID outperforms in most metrics, except for slightly higher occlusion compared to Pipeline 1, which places the product centrally with a small bounding box.
+T-Stars-Poster outperforms in most metrics, except for slightly higher occlusion compared to Pipeline 1, which places the product centrally with a small bounding box.
 However, Pipeline 1 uses GPT-4o for layout, lacking product-centric design knowledge, resulting in lower quality.
-PAID uses trained experts for different tasks in ad design, creating more visually appealing results (see Figure 5).
-Pipeline 2 fixes the product location, limiting adaptability to product shapes, whereas PAID adjusts layouts based on product shape for better visuals.
-Additionally, the fixed background of Pipeline 2 can lead to layouts with crowded tagline boxes, making them hard to read, while our model provides more flexible and readable layouts.
+T-Stars-Poster uses trained experts for different tasks in ad design, creating more visually appealing results (see Figure 5).
+Pipeline 2 fixes the product location, limiting adaptability to product shapes, whereas T-Stars-Poster adjusts layouts based on product shape for better visuals.
+Additionally, the fixed background of Pipeline 2 can crowd tagline boxes, reducing readability, while our model generates more flexible and readable layouts.
 Table 1. Quantitative comparison between automatic pipelines for product-centric advertising image design.
@@
 Pipeline 2 43.221 1.814 0.0016 0.0028 0.9994 0.9930 0.0984 0.1158 0.1864
-PAID 37.524 1.405 0.0013 0.0017 0.9999 0.9973 0.1367 0.0955 0.1815
+T-Stars-Poster 37.524 1.405 0.0013 0.0017 0.9999 0.9973 0.1367 0.0955 0.1815
 Compare with prompt generation methods.
@@
 The FBM Rate is determined by votes from three annotators for each case.
-As shown in Table 2, our method outperforms the baselines in both FBM Rate and ED Score.
-This means that our model effectively aligns with the e-commerce prompt dataset distribution after training.
-We also visualize the feature distribution of prompts generated in the appendix.
+As shown in Table 2, our method outperforms the baselines, which means that our model effectively aligns with the e-commerce prompt dataset distribution after training.
 Table 2. Evaluation of prompt generation methods.
@@
 Compare with layout generation methods.
-To verify the advantage of our layout generation model, we compare it with previous background-based methods (DS-GAN (Hsu et al., 2023), RADM (Li et al., 2023b), and PosterLlama (Seol et al., 2024)) and the foreground-based method P&R (Li et al., 2023a).
-For background-based generation methods, we conduct experiments on two versions of the testing set for fair comparison, named the erased/generated set.
-The erased set contains images with graphic elements erased by LaMa (Suvorov et al., 2022), which may provide hints to put graphic elements on these erased areas.
-The generated set contains images re-generated using their ground-truth prompts (image captions).
-The generated set has no above hints and conforms to real applications.
-As for the foreground-based method P&R, since it has not released the codes, we re-implement its core idea: first, generating the layout of the foreground and graphic elements conditioned on the foreground and taglines, and second, generating the final image based on the background prompt, layout, and foreground.
-Our method uses the ground-truth prompt, product foreground, and taglines for prediction.
+We compare our methods with background-based methods (DS-GAN (Hsu et al., 2023), RADM (Li et al., 2023b), and PosterLlama (Seol et al., 2024)) and the foreground-based method, P&R (Li et al., 2023a).
+For background-based methods, we test on two sets for fair comparison: the erased set, with graphic elements erased by LaMa (Suvorov et al., 2022), which may provide hints to put graphic elements on erased areas; and the generated set, which contains re-generated images with their ground-truth prompts (image captions), aligning with real applications.
+Since P&R has not released the codes, we re-implemented its idea.
+Ground-truth prompts are also given to P&R and our method.
 The results are shown in Table 3 and Figure 6.
-Background-based methods perform relatively poorly in terms of product occlusion and overlap and alignment of graphic elements.
-This may be due to the fixed background image content, leaving less room for the model flexibility.
-RADM and DS-GAN, which are insensitive to tagline content, also face issues with generating tagline boxes that cannot adequately fill the tagline content.
-The foreground-based method P&R improves in these areas but does not consider the harmony between the product, background prompt, and graphic layout, sometimes leading to unreasonable generated images.
-Our method, PAID, takes into account the shape of the product foreground, background prompt, and tagline content to decide the overall layout, including graphic and nongraphic elements, resulting in a better overall effect.
+Background-based methods struggle with product occlusion, overlap, and alignment of graphic elements, likely due to fixed background content.
+RADM and DS-GAN also have issues with tagline boxes that cannot properly fit the content.
+P&R improves in these areas but sometimes generates unreasonable images due to a lack of harmony between product, background prompt, and graphic layout.
+T-Stars-Poster considers product shape, background prompt, and tagline content, optimizing the entire layout for graphic and non-graphic elements, ultimately achieving better overall results.
 Table 3.
@@
 Compare with layout-controlled inpainting models.
-We compare our layout-controlled background image generation model with others, including training-based method GLIGEN (Li et al., 2023c) and training-free methods BoxDiff (Xie et al., 2023), TFCLG (Chen et al., 2024), and MultiDiffusion (Bar-Tal et al., 2023).
+We compare our layout-controlled background image generation model with others: the training-based method GLIGEN (Li et al., 2023c) and training-free methods like BoxDiff (Xie et al., 2023), TFCLG (Chen et al., 2024), and MultiDiffusion (Bar-Tal et al., 2023).
 Since these methods lack inpainting ability, we re-implement them with the SDXL-based inpainting ControlNet (Creative, 2024).
-Quantitative results are reported in Table 4.
-Training-free methods have weaker spatial control compared to training-based methods.
+The results are reported in Table 4.
 Our trained model outperforms others in FID, mIoU, and AP, indicating better image quality and spatial control.
-As shown in Figure 7, other models tend to miss some instances when more objects are required, while our model generates them more accurately.
-Figure 7. Visualization of images generated by different layout-controlled inpainting models.
-Table 4. Quantitative performance comparison of layout-controlled inpainting models on PIL test set.
+Table 4. Quantitative comparison of layout-controlled inpainting models on PIL test set.
 […]
@@
 Ours 25.917 0.313 0.906 0.705 0.079/0.127/0.078
+5.4. User-Centric Study
+To thoroughly evaluate the visual effects and attractiveness of the generated results, we randomly select 200 test samples and ask 7 Taobao users to compare images in pairs, evaluating background aesthetics, layout harmony, text readability, and overall visual appeal.
+As shown in Table 5, our method significantly outperforms the comparison methods in all aspects.
+Table 5. User-centric study. The values in the table represent the winning rate of our method compared to other methods.
+Comparison Method Bg Aesthetic Layout Harmony Text Readability Visual Appeal
+Pipeline 1 84% 93% 94% 91%
+Pipeline 2 77% 85% 89% 74%
 5.5. Online Result
-To assess the online performance of PAID, we conduct A/B tests in two advertising recommendation scenarios on Taobao.
-We randomly select 5,000 products from various categories (including clothing, cosmetics, food, and electronics) and collect their foreground images and taglines from the advertisers.
-We then generate advertising images for each product using both PAID and Pipeline 2 (the top-performing baseline as shown in Table 1).
-The A/B tests are conducted using 5% of the main traffic, affecting only the experimental products when this traffic is directed to them.
-After gathering data over one month, we found that PAID achieved a 3.02% and 3.03% increase in click-through rate (CTR) for the two recommendation scenarios, respectively.
-This indicates that PAID is capable of producing more visually appealing advertising images, resulting in improved recommendation outcomes.
+To evaluate the online performance of T-Stars-Poster, we conduct A/B tests in two advertising recommendation scenarios on Taobao.
+We randomly select 5,000 products and collect their foreground images and taglines from advertisers.
+We generate images for each product using both T-Stars-Poster and Pipeline 2 (the top-performing baseline as shown in Table 1).
+The A/B tests use 5% of the main traffic, affecting only the experimental products when this traffic is directed to them.
+After gathering data over one month, results show that T-Stars-Poster achieved a 3.02% and 3.03% increase in click-through rate (CTR) in the two recommendation scenarios, respectively.
+This indicates that T-Stars-Poster produces more visually appealing advertising images, resulting in improved recommendation outcomes.
+6. Limitations and Discussion
+Through observation, we have found that the results may be unsatisfactory in some cases.
+1) Difficulty in adjusting product angle and lighting.
+As shown in Figure 5, inpainting models cannot adjust the product angle and lighting due to the model structure and training data, possibly causing disharmony in lighting and composition.
+While subject-driven methods (Tan et al., 2024) may help, they currently cannot fully preserve product features.
+2) Limited flexibility with multiple foregrounds.
+As Figure 5 shows, multiple foregrounds are treated as one entity.
+This limits adjustments when products are spaced apart or angled very differently.
+Future improvements could involve treating separate subjects individually and introducing position augmentations during training.
+3) Poor layout with many taglines.
+Over five taglines increase overlap and affect readability due to limited training data (4.3%) and canvas space.
+More data with multiple taglines might help resolve this.
 7. Conclusion
-In this paper, a novel framework named Product-Centric Advertising Image Design (PAID) is proposed to automatically create advertising images only based on product foreground images, marketing taglines, and target sizes.
-PAID consists of four stages: prompt generation, layout generation, background image generation, and graphics rendering.
-First, it generates adaptive prompts according to the product foregrounds.
+This paper introduces T-Stars-Poster, a product-centric framework for creating advertising images using product foregrounds, taglines, and target sizes.
+It comprises four stages: prompt generation, layout generation, background image generation, and graphics rendering.
+First, it generates prompts according to the product foregrounds.
 Then, it predicts how graphic and nongraphic elements should be placed according to the prompt, product foreground, and taglines, creating a harmonious overall layout.
 Next, a layout-controlled inpainting model is utilized for background image generation.
-A graphics rendering module is then applied to get the final images.
+A graphics rendering module is applied to get the final images.
 We train separate experts to conduct these sub-tasks.
 Two datasets are created for the convenience of training and testing.
-The results of the test set and the online A/B tests have proven that the proposed PAID framework generates more visually pleasing and attractive advertising images.
+Test results and online A/B tests show that T-Stars-Poster produces more visually pleasing and attractive advertising images.
 Acknowledgements
-This work is supported by National Natural Science Foundation of China (NSFC No.62076031 and No.62076036) and Alibaba Research Intern Program.
+This work is supported by Alibaba Research Intern Program.
 Appendix A Dataset Examples
-In this section, we provide examples from the PITA and PIL datasets, illustrated in Figures 8 and 9, respectively.
+In this section, we provide examples from the PITA and PIL datasets, illustrated in Figures 7 and 8, respectively.
 The PITA dataset includes advertising images with marketing taglines.